### v1

# Light Scalars in Light of UV/IR Mixing Classicalization via Synergy between Vainshtein & Chameleon Screenings

Abstract
Effective field theories featuring light scalar fields play a pivotal role in addressing fundamental questions in (astro)particle physics and cosmology. However, such theories often confront hierarchy problems in the absence of a symmetry. Self-completion via classicalization offers a non-Wilsonian approach to ultraviolet (UV) completion, wherein new scalar self-interactions involving derivatives give rise to Vainshtein-like screening around energy-momentum sources. Rather than introducing new UV degrees of freedom to restore unitarity at high energies, these theories reshuffle their infrared (IR) degrees of freedom by generating extended semi-classical objects—referred to as classicalons—which decay into a multitude of soft particles. This mechanism incorporates non-localizable fields, thereby realizing a form of UV/IR mixing that is analogous to the dynamics of black holes in gravitational theories. In this article, having reviewed the fundamental principles of classicalization with a simple k-essence model, we then argue the necessity of maintaining a little hierarchy between the scalar mass and the scale of the first new resonances, thereby illustrating the impact of UV/IR mixing on hierarchy problems. Additionally, we investigate the effects of a scalar potential and couplings to fermions on the Vainshtein screening mechanism. We discuss that a chameleon-like screening mechanism must accompany the Vainshtein screening to preserve the integrity of classicalon solutions.
1 Introduction
Relatively light scalar fields are ubiquitous in theoretical (astro)particle physics and cosmology. They serve as essential components in numerous models, including those involving spontaneous symmetry breaking (SSB) [1], particle dark matter [2], dynamical dark energy [3], and inflation [4], to cite only a few representative reviews from the literature.
From an effective field theory (EFT) perspective [5], light scalars face hierarchy problems [6]: their masses require protection from radiative corrections induced by heavier particles. The principle of ’t Hooft naturalness [7] traditionally suggests that a symmetry should underlie their lightness, an approach that remains the most widely adopted. However, this symmetry-based explanation has been called into question by the discovery of the Higgs boson at the CERN LHC and ““nothing else”” [8, 9, 10], motivating the search for alternative mechanisms to account for the lightness of scalar fields [11, 12]. In recent years, promising new approaches have emerged, including cosmological selection mechanisms (see Ref. [13] for a review), and the so-called ‘accidents’ [14, 15]. In this article, we explore a more radical possibility inspired by quantum gravity considerations: ultraviolet (UV)/infrared (IR) mixing.
The hierarchy problems associated with weakly coupled scalars originate from the Wilsonian perspective of quantum field theories (QFTs) as EFTs [16], where the decoupling of scales is of paramount importance. It is crucial to recognize that the principles of decoupling and the axioms of textbook QFT [17] are deeply intertwined. Nevertheless, it has been established that the locality axiom, typically enforced through microcausality [17], is overly restrictive when addressing non-perturbative phenomena in gravity [18, 19, 20, 21, 22, 23, 24, 25], such as black holes (BHs). This suggests that the realization of locality in the fundamental theory of nature may differ substantially from conventional wisdom. If a certain degree of non-locality is permitted within QFT, one can envision correlations between short- and long-range physics—referred to as UV/IR mixing—which could explain the apparent violation of naturalness in the EFT from a naive Wilsonian perspective [26, 27, 28, 29, 30, 31, 32, 33, 34, 11, 12, 35].
[…]
$\Lambda_{P}\equiv 1/\ell_{P}\equiv\sqrt{\frac{1}{8\pi G_{N}}}\sim 10^{18}\ \text{GeV},$ (1.1)
at which non-perturbative effects of quantum gravity become significant. Consequently, strong-field gravity inverts the conventional relationship between energy-momentum and spacetime resolution, as depicted in Fig. 1.
Figure 1: Schematic representation of UV/IR mixing in gravity upon crossing the Planck scale, $\Lambda_{P}$. A pointlike object with mass $M\ll\Lambda_{P}$ is accurately described as a quantum particle, for which pair creation becomes significant at distances below the quantum wavelength $\Delta L\sim 1/M$ (with gravitational effects remaining negligible). Conversely, for $M\gg\Lambda_{P}$, the object is most appropriately characterized as a classical black hole, where gravitational effects dominate within the gravitational radius $\Delta L\sim\ell_{P}^{2}M$ (with quantum fluctuations being negligible). At the boundary where $M\sim\Lambda_{P}$, the length scale satisfies $\Delta L\sim\ell_{P}$, and both quantum fluctuations and gravitational effects become equally significant. The theoretical description of such a state remains an open question. The rainbow background in the figure highlights this UV/IR mixing, with the inversion of the relationship between mass and length scale when crossing $\Lambda_{P}$.
This inversion leads to the hypothesis known as ‘asymptotic darkness’ [39, 40, 41]: an ultra-Planckian scattering process between 2 particles, with a center-of-mass energy $\sqrt{s}\gg\Lambda_{P}$ and an impact parameter $b\lesssim R_{S}\sim\ell_{P}^{2}\sqrt{s}$, should result in the formation of a BH of mass $M\sim\sqrt{s}$. This BH subsequently evaporates—via Hawking radiation [42, 43]—into $N_{\circledast}\sim R_{S}\sqrt{s}\gg 1$ soft particles, each with energy $\omega\sim 1/R_{S}$. This scenario is supported by the $S$-matrix program in gravity [24, 25], both in GR and string theory [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. As a result, $\ell_{P}$ emerges as the smallest length scale accessible in any experiment, rendering the introduction of new degrees of freedom with mass $M_{\text{new}}\gg\Lambda_{P}$ meaningless, as they cannot be distinguished from a BH of equivalent mass [40, 41]. It follows that enforcing locality as a strict microcausality condition lacks physical meaning in a theory characterized by a minimal length scale [61], where light cones inherently appear ‘fuzzy’ [62, 63].
A corollary of asymptotic darkness is that pure gravity in GR, when treated as a QFT [64, 65, 66], may be self-complete in the ultra-Planckian regime [40, 41, 67]. The transition amplitude for $2\to 2$ ultra-Planckian hard scattering processes of gravitons—processes that violate perturbative unitarity—is suppressed by a factor of $e^{-N_{\circledast}}$, since a BH-like object is produced and preferentially decays into $N_{\circledast}\gg 1$ soft particles [67, 59, 60]; unitarity is thus restored non-perturbatively. To reconcile a quantum description of BHs with unitarity, a natural framework—proposed in Refs. [68, 69, 70, 71, 72, 73]—is to interpret them as coherent states composed of $N_{\circledast}\gg 1$ gravitons [^2] For comprehensive reviews and additional references, see Refs. [74, 75]. Subsequent developments are discussed in Refs. [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. Although not widely adopted as a quantum description of BHs, this framework simply follows the textbook treatment of a semi-classical background for a bosonic field in QFT: a coherent state with a large occupation number [17].. This corpuscular description of gravity has also been extended to other gravitational backgrounds of cosmological significance, yielding important implications for the cosmological constant problem [87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97].
Should new states be necessary to restore unitarity at energies $\sqrt{s}\sim\Lambda_{P}$, they must consequently emerge at or below the scale $\Lambda_{P}$, where gravitational interactions are expected to become strongly coupled. Around the Planck scale, new resonances—sometimes dubbed quantum BHs [98, 99, 100, 101], Planckions [102, 103], or BH precursors [104, 105]—should appear, the properties of which demand a UV-complete theory of quantum gravity for their description. One may then hypothesize that these new states could represent composite resonances of gravitons, potentially forming their own composite gravitational strings before transitioning into the ultra-Planckian BH regime [40]. This scenario is analogous to the regulation of IR behavior in quantum chromodynamics (QCD), where the strong interaction between quarks and gluons generates hadronic strings [106]. While this hypothesis is not logically inconsistent, its investigation requires non-perturbative techniques in quantum gravity, such as lattice approaches [^3] The modern approach to lattice gravity is known as causal dynamical triangulations (CDT) [107]..
Inspired by the foregoing considerations, the concept of ‘classicalization’—first introduced in Ref. [108] and subsequently developed in Refs. [109, 110, 111, 67, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 103, 125, 79, 126]—posits that certain EFTs may not admit a conventional Wilsonian UV-completion. Instead, they might achieve self-UV-completion [^4] The ‘self-healing’ mechanism, as discussed in Ref. [127], reveals that the scale at which tree-level unitarity is violated does not necessarily align with the emergence of new physics. This mechanism resolves the apparent unitarity violation within the EFT framework, obviating the need for additional degrees of freedom. by unitarizing through the non-perturbative formation of extended semi-classical objects termed ‘classicalons’ (analogous to BHs in gravity). This mechanism involves transitioning through a strongly coupled intermediate regime of composite states—referred to as ‘fuzzyons [^5] The term ‘fuzzyons’ alludes to the minimal length scale $\ell_{\ast}$ at which the classicalizer field makes new types of composite states.’ in Ref. [128]—at a characteristic scale $\Lambda_{\ast}\equiv 1/\ell_{\ast}$, and the quasi-continuum of classicalon states emerges in the deep-UV region of the spectrum [112] (see Fig. 2). Thus, UV/IR mixing constitutes the ‘DNA’ of classicalization. Classicalizing theories have been identified as non-localizable [^6] Based on the construction in Ref. [129], where non-local form factors appear even at tree level (unlike in the present work), the model of Ref. [128] has been argued to exhibit classicalization. For alternative constructions of this type with a classicalizing behavior, see also Refs. [130, 131, 76, 132]. QFTs, in which microcausality is superseded by macrocausality [133, 134, 135].
The remarkable phenomenon of classicalization is not exclusive to gravitons. For instance, in modified gravity, it is common to encounter additional forces mediated by new scalar fields with derivative self-couplings, which have significant implications for cosmology (see Ref. [3] for a comprehensive review). These models exhibit the Vainshtein [^7] The literature on dark energy theories lacks a standardized nomenclature for screening mechanisms. In this article, we adopt the terminology of Ref. [136], wherein mechanisms relying on non-linearities in the field derivatives are referred to as Vainshtein-like screening, and those based on non-linearities in the fields themselves are termed chameleon-like screening. For brevity, we omit the suffix “like”. screening mechanism [137] around astrophysical bodies, enabling them to remain consistent with the stringent constraints on new long-range forces within the solar system while still being active at cosmological scales. Notable examples include k-essence models [138, 139, 140, 141, 142] and Galileon theories [143, 144, 145, 146]. Vainshtein screening appears fundamentally incompatible with a Wilsonian UV-completion [147, 148, 149, 150], and classicalization remains the only known possibility for a UV-completion [^8] Unfortunately, the large length scales involved in dark energy models led the authors of Ref. [151] to conclude that such scenarios should be excluded by cross-section measurements at hadron colliders. This criticism does not pertain to other applications featuring a sufficiently large classicalizing energy scale. [152, 116, 133, 134, 153, 151]. This outcome is not unexpected, as Vainshtein screening inherently exhibits UV/IR mixing [133, 134].
Figure 2: Panel $(a)$: Classicalization is governed by an interaction scale $\Lambda_{\ast}$. In a hard scattering process with center-of-mass energy $\sqrt{s}$, 3 distinct regimes can be identified: $(i)$ the production of $\mathcal{O}(1)$ weakly interacting bosons when $\sqrt{s}\ll\Lambda_{\ast}$; $(ii)$ the production of a narrow, strongly coupled resonance (a fuzzyon) decaying into $N_{\circledast}\sim 1$ bosons when $\sqrt{s}\sim\Lambda_{\ast}$; $(iii)$ the production of a semi-classical state (a classicalon) decaying into $N_{\circledast}\gg 1$ soft bosons when $\sqrt{s}\gg\Lambda_{\ast}$. Panel $(b)$: Schematic representation of the mass spectrum of composite states of $N_{\circledast}$ bosons in a theory exhibiting classicalization (logarithmic scale). As discussed in Ref. [154, 112], this spectrum is expected to be quantized as a function of the interaction scale $\Lambda_{\ast}$ and a real parameter $\gamma>1$, which depends on the operator responsible for triggering classicalization. In a collider experiment with a center-of-mass energy $\sqrt{s}\sim\Lambda_{\ast}$, only quantum composite states (fuzzyons) within the strongly coupled regime can be probed $(N_{\circledast}\sim 1)$. For $\sqrt{s}\gg\Lambda_{\ast}$, however, the states (classicalons) form a quasi-continuum that is effectively described by a semi-classical approach $(N_{\circledast}\gg 1)$.
In this article, we examine the field-theoretic aspects of classicalization for gauge singlet scalars, independent of its specific applications in (astro)particle physics or cosmology. In Section 2, we review the interplay between Vainshtein screening and classicalization within a unified framework, adopting the perspective of self-UV-completion rather than the conventional motivation of screening scalar forces in dark energy theories. We also summarize the criticisms in the literature concerning classicalization, along with their counterarguments, and elucidate the fundamental distinction between the standard EFT interpretation of non-renormalizable interactions and their interpretation within the framework of classicalization. In Section 3, we investigate the constraints imposed by introducing a potential that does not disrupt the screening mechanism—a consideration overlooked in previous studies. In particular, we provide a detailed analysis of the claim in Ref. [108] that the mediator must be lighter than the classicalizing scale $\Lambda_{\ast}$, thereby suggesting the existence of light scalars with a necessary but little hierarchy. Then, we explore, for the first time, the consequences of coupling the scalar field to other fields, with a focus on fermions. In particular, we demonstrate the necessity of incorporating a variant of the chameleon screening mechanism—well-known in dark energy theories [3]—to preserve an active Vainshtein screening within the theory when a scalar potential and/or a direct coupling to matter are present. In Section 4, we summarize our results and discuss potential avenues for future research. N.B.: Our conventions follow those adopted in the QFT textbook by Peskin and Schroeder [1].
2 K-essence as Classicalizer
For pedagogical purposes, we begin with a brief review of the classicalization proposal. We then proceed to examine a k-essence model involving a massless scalar field characterized by a simple kinetic self-interaction. This section establishes the foundational results upon which the original analysis in Section 3 is based.
2.1 Classicalization in a Nutshell
The classicalization phenomenon [108] relies on a bosonic ‘classicalizer’ field that generalizes the role of gravity by generating extended classical field configurations, known as classicalons. For this mechanism to operate, the classicalizer must couple to an operator that becomes strongly interacting at short distances—such as in a scattering process—thereby inducing the formation of a classicalon. When derived exclusively from self-sourcing—that is, in the absence of external sources—classicalons are non-topological solitons [155] belonging to the universal class of objects known as ‘saturons,’ whose defining characteristic is the saturation of unitarity and entropy bounds [156, 157, 158, 159, 80, 160, 82, 161].
To illustrate this concept, consider a scalar classicalizer field $\phi(x)$ of mass $m$ defined on a $3+1$-dimensional Minkowski spacetime $\mathbb{R}^{1,3}$. A prototypical classicalizing operator—referred to as ‘UV-screener’—takes the form:
$\frac{\phi}{\Lambda_{\ast}^{d-4}}\,\mathcal{O}_{S}^{(d-1)},$ (2.1)
where $\mathcal{O}_{S}^{(d-1)}$ is a scalar composite operator [^9] One could also consider an operator of the form: $\frac{\partial_{\mu}\phi}{\Lambda_{\ast}^{d-4}}\,\mathcal{O}_{V}^{\mu\,(d-2)},$ (2.2) in which the classicalizer field couples to a vector operator $\mathcal{O}_{V}^{\mu\,(d-2)}$ of dimension d2d-2. However, this operator is equivalent to the scalar case in Eq. (2.1), since integration by parts yields $\int d^{4}x\ \frac{\partial_{\mu}\phi}{\Lambda_{\ast}^{d-4}}\,\mathcal{O}_{V}^{\mu\,(d-2)}=\int d^{4}x\ \frac{\phi}{\Lambda_{\ast}^{d-4}}\,\mathcal{O}_{S}^{(d-1)},\quad\text{with}\quad\mathcal{O}_{S}^{(d-1)}=-\partial_{\mu}\mathcal{O}_{V}^{\mu\,(d-2)},$ (2.3) assuming the boundary term vanishes. of dimension $d-1\geq 3$. This operator encodes a non-renormalizable interaction involving $\phi(x)$ and/or other fields, with the interaction scale governed by $\Lambda_{\ast}\equiv 1/\ell_{\ast}$.
From a conventional Wilsonian perspective, the Lagrangian density resembles that of an EFT, where the perturbative $\phi\phi\to\phi\phi$ scattering amplitude $\mathcal{M}(s,t)$—expressed in terms of the Mandelstam variables $s$ and tt—violates unitarity in high-energy scattering processes when $s\sim-t\gtrsim\Lambda_{\ast}^{2}$. Traditionally, this violation would necessitate the introduction of new degrees of freedom to restore theoretical consistency. Classicalization, however, offers a distinct, non-Wilsonian UV-completion where unitarity is restored non-perturbatively [108, 103]. Specifically, the formation of a classicalon of mass $M_{\circledast}\sim\sqrt{s}$ and radius
$R_{\circledast}=\ell_{\ast}\left(\frac{M_{\circledast}}{4\pi\Lambda_{\ast}}\right)^{\gamma-1},\quad\gamma>1,$ (2.4)
which grows with a model-dependent exponent $\gamma-1$ of the center-of-mass energy $\sqrt{s}\gg\Lambda_{\ast}$, prevents the localization of $\phi$-particles within a distance shorter than $R_{\circledast}(s)\gg\ell_{\ast}$. This classicalon can be described as a coherent state [114, 119, 121, 126] of
$N_{\circledast}=\frac{M_{\circledast}}{\omega}\gg 1$ (2.5)
weakly interacting $\phi$-bosons confined within a region of size $2R_{\circledast}$. The energy of each constituent boson is typically [^10] A Bose-Einstein statistical model of the classicalon is developed in Ref. [114], and its results align with the semi-classical coherent-state approach described in Ref. [119].:
$\omega=\sqrt{\left(\frac{\pi}{2R_{\circledast}}\right)^{2}+m^{2}}>m.$ (2.6)
Like BHs undergoing Hawking radiation, the classicalon evaporates thermally into $N_{\circledast}\gg 1$ quanta.
Starting from an $\mathcal{O}(1)$ number of bosons in the initial state, the amplitude for producing a specific $N_{\circledast}$-boson microstate of the classicalon $\mathfrak{S}$ is exponentially suppressed as $\sim e^{-N_{\circledast}}$. However, from the perspective of the $S$-matrix, individual microstates are indistinguishable. Consequently, one must perform an inclusive sum over the vast number of $\sim e^{+N_{\circledast}}$ microstates of $\mathfrak{S}$, which compensates for the exponential suppression [108, 67, 114, 119, 158, 126]. Unitarity is preserved in the $2\to 2$ process because it proceeds via an intermediate classicalon state, $2\to\mathfrak{S}\to 2$. The probability for any single microstate to decay back into 2 particles is suppressed as $\sim e^{-N_{\circledast}}$.
The core idea of classicalization is to transform a $2\to 2$ hard scattering process—one that would otherwise violate perturbative unitarity—into a soft $2\to N_{\circledast}\gg 1$ process, thereby realizing UV/IR mixing. This can be interpreted as a collection of feeble elementary interactions among the $N_{\circledast}$ classicalon’s constituents. If these bosons are massive, their average individual energy $\omega>m$ must be hierarchically smaller than the interaction scale $\Lambda_{\ast}$; otherwise, the $N_{\circledast}$ bosons would not reside in the feebly interacting regime of the EFT. This requirement explains why UV/IR mixing necessitates a little hierarchy $m\ll\Lambda_{\ast}$, as previously suggested in Ref. [108].
Figure 3: Feynman diagram of a non-perturbative process, $2\to N_{\circledast}\gg 1$, mediated by the production and evaporation of a classicalon involving a k-essence field $\phi$.
2.2 Classicalization by Vainshtein Screening
2.2.1 Prototype Model
Consider the following Lagrangian density $\mathcal{L}_{X}$ for a real scalar field $\phi(x)$ of mass dimension 1 (the classicalizer), defined on $\mathbb{R}^{1,3}$:
$\mathcal{L}_{X}\equiv\Lambda_{\ast}^{4}\,\mathcal{K}(X),\quad\text{where}\quad X\equiv\dfrac{\partial^{\mu}\phi\partial_{\mu}\phi}{2\Lambda_{\ast}^{4}}.$ (2.7)
This Lagrangian density depends solely on the kinetic variable $X$. For simplicity, we adopt the following kinetic function:
$\mathcal{K}(X)=X+c_{2}X^{2},\quad\text{with}\quad c_{2}\equiv\pm 1,$ (2.8)
which yields the Lagrangian density,
$\mathcal{L}_{X}=\dfrac{1}{2}\,\partial^{\mu}\phi\partial_{\mu}\phi+\dfrac{c_{2}}{4\Lambda_{\ast}^{4}}\left(\partial^{\mu}\phi\partial_{\mu}\phi\right)^{2},$ (2.9)
of a prototype model of k-essence. Here, the second term represents a non-renormalizable kinetic self-interaction. The Lagrangian density $\mathcal{L}_{X}$ exhibits both a shift symmetry $\phi(x)\mapsto\phi(x)+\phi_{c}$ (where $\phi_{c}\in\mathbb{R}$ is a constant) and a $\mathbb{Z}_{2}$ symmetry $\phi(x)\mapsto-\phi(x)$.
At first glance, this theory is formulated as an EFT, truncated at the leading self-interaction term permitted by these symmetries, with a naive perturbative cutoff $\Lambda_{\ast}\equiv 1/\ell_{\ast}$. However, if this theory self-completes through classicalization, it implies that it is secretly UV-complete, despite its initial appearance as an EFT. Our objective is to review the plausibility of this assertion.
To determine whether this theory leads to classicalization [^11] In the original article [108], the authors also considered the classicalization of theories described by the action: $\int d^{4}x\ F^{2}(\phi)\,\partial^{\mu}\phi\partial_{\mu}\phi,$ (2.10) where $F(\phi)$ is a function of $\phi$ satisfying $F^{2}(0)=1/2$. However, as discussed in Ref. [162], such a Lagrangian can be reformulated as the kinetic term of a free theory for a field $\widetilde{\phi}$ through the field redefinition $\widetilde{\phi}(\phi)$, defined by $d\widetilde{\phi}/d\phi\equiv F(\phi)$. Therefore, such a theory is secretly a non-interacting theory and does not exhibit classicalization., we must verify the following conditions: $(i)$ the existence of a classicalon solution in the presence of a sharply localized source; $(ii)$ the stability of this classicalon solution against quantum corrections; and $(iii)$ whether the non-linearities arising from the kinetic self-coupling prevent the localization of $\phi$-quanta within the classicalization radius.
2.2.2 Classicalon Solution from Classical Vainshtein Screening
To determine whether the kinetic self-interaction in $\mathcal{L}_{X}$ can induce classicalization, we must first verify the existence of a classicalon solution within a semi-classical framework. To enable an analytic derivation of this solution, we adopt the approach outlined in Ref. [108] and investigate the response of $\phi(x)$ to an external, localized energy-momentum source $J(x)$. The corresponding term introduced into the Lagrangian is given by
$\mathcal{L}_{J}=\frac{\phi}{\Lambda_{\ast}}\,J,\quad\text{where}\quad J=-\mathcal{E}\,\delta^{(3)}(\mathbf{r}),\quad\text{and}\quad\int d^{3}\mathbf{r}\ \delta^{(3)}(\mathbf{r})=1.$ (2.11)
Here, $J$ represents a pointlike source [^12] In dark energy theories, the magnitude $\mathcal{E}$ of the localized source is represented by the mass of the astrophysical object (e.g., a star) around which Vainshtein screening takes place [3]. Naturally, the pointlike approximation should be relaxed when required. of magnitude $\mathcal{E}>0$. The underlying concept is that if a classicalon can be generated by a pointlike external source, it will necessarily also arise from the kinetic self-interaction when attempting to sharply localize a wave packet of $\phi$-bosons. It is important to note that, while the source term $\mathcal{L}_{J}$ preserves shift symmetry, it explicitly breaks Lorentz-Poincaré invariance and the $\mathbb{Z}_{2}$ symmetry.
The first step in a semi-classical analysis is to derive the classical background solution for the field. The Euler-Lagrange equation for the Lagrangian density $\mathcal{L}_{X}+\mathcal{L}_{J}$ is given by
$\square\phi+\dfrac{c_{2}}{\Lambda_{\ast}^{4}}\,\partial^{\mu}\left(\partial^{\nu}\phi\partial_{\nu}\phi\partial_{\mu}\phi\right)=-\dfrac{\mathcal{E}}{\Lambda_{\ast}}\,\delta^{(3)}(\mathbf{r}),$ (2.12)
which, for a static source, simplifies to
$\overrightarrow{\nabla}\cdot\left[\overrightarrow{\nabla}\phi-\dfrac{c_{2}}{\Lambda_{\ast}^{4}}\left(\overrightarrow{\nabla}\phi\right)^{2}\overrightarrow{\nabla}\phi\right]=\dfrac{\mathcal{E}}{\Lambda_{\ast}}\,\delta^{(3)}(\mathbf{r}).$ (2.13)
Given the spherical symmetry of the source, we adopt spherical coordinates $\mathbf{r}\equiv(r,\theta,\varphi)$. Applying the divergence theorem [^13] Also referred to as Gauss’s theorem or the Gauss-Ostrogradsky theorem., we integrate the above equation to obtain:
$\phi^{\prime}-\dfrac{c_{2}}{\Lambda_{\ast}^{4}}\,\phi^{\prime\,3}=\frac{1}{\Omega r^{2}}\cdot\dfrac{\mathcal{E}}{\Lambda_{\ast}},$ (2.14)
where $\phi^{\prime}\equiv d\phi/dr$, and $\Omega\equiv 4\pi$ denotes the solid angle. This cubic equation in $\phi^{\prime}$ admits a solution that, while not particularly illuminating in its full form [3], allows us to identify 3 distinct regimes: $(i)$ the linear regime, characterized by the hierarchy $\phi^{\prime}\gg\ell_{\ast}^{4}\,\phi^{\prime\,3}$; $(ii)$ the non-linear regime, characterized by the opposite hierarchy $\phi^{\prime}\ll\ell_{\ast}^{4}\,\phi^{\prime\,3}$; and $(iii)$ the transition regime, marking the onset of non-linearities at the length scale:
$R_{V}\equiv\ell_{\ast}\sqrt{\frac{\mathcal{E}}{\Omega\Lambda_{\ast}}},$ (2.15)
known as the Vainshtein radius.
The background solution $\overline{\phi}^{\,\prime}(r)$ in these regimes is expressed as
[…]
where a real solution in the non-linear regime exists only if $c_{2}=-1$. Notably, for all regimes,
$\left|\dfrac{\overline{\phi}(r)}{r\,\overline{\phi}^{\,\prime}(r)}\right|=\mathcal{O}(1).$ (2.17)
Here, the non-linearities implement the classical version of the Vainshtein screening mechanism: the ratio of the background solutions $\overline{\phi}^{\,\prime}(r)$, with and without the kinetic self-interaction term, scales as $(r/R_{V})^{4/3}\ll 1$ at distances $r\ll R_{V}$ from the source at $r=0$. Consequently, when non-linearities dominate, they suppress the scalar force in the region surrounding the source, see Fig. 4.
Figure 4: Comparison of the exact background solution $\overline{\phi}^{\prime}(r)$ with the asymptotic solutions in both the linear and non-linear regimes, demarcated by the Vainshtein radius $R_{V}$, at which $\overline{\phi}^{\prime}(r)\sim\Lambda_{\ast}^{2}$, the classicalizing scale.
The preceding discussion yields several key observations:
  • $\spadesuit$
    The Vainshtein radius, defined in Eq. (2.15), represents the transition scale between the linear and non-linear regimes, growing with $\mathcal{E}$. For $\mathcal{E}\ll\Lambda_{\ast}$, the radius satisfies $R_{V}\ll\ell_{\ast}$, and non-linearities remain negligible within the valid regime of the EFT. In contrast, for $\mathcal{E}\gg\Lambda_{\ast}$, the radius becomes $R_{V}\gg\ell_{\ast}$, and the source is enclosed by a spherical region of radius $R_{V}$ where non-linearities dominate (the ‘Vainshtein core’). A notable puzzle in this k-essence model is that, although $R_{V}\gg\ell_{\ast}$ lies within the EFT’s valid length scale, the condition $\phi^{\prime}(r\ll R_{V})\gg\Lambda_{\ast}$ appears to exceed the regime of validity in terms of the background amplitude $\overline{\phi}^{\,\prime}(r)$ [134]. Assessing the radiative stability of this background is therefore of paramount importance to validate the reliability of this semi-classical analysis.
  • $\spadesuit$
    The requirement $c_{2}=-1$ violates positivity bounds [163, 147, 116, 164, 165]—a point to which we will return in Section 2.3.3—whereas $c_{2}=+1$ guarantees the existence of a Wilsonian UV-completion. This suggests that Vainshtein screening occurs only when the UV-completion is non-Wilsonian.
Taken together, these characteristics serve as signatures of the UV/IR mixing induced by the response of the derivative self-sourcing term to a localized source with $\mathcal{E}\gg\Lambda_{\ast}$.
2.2.3 Radiative Stability from Quantum Vainshtein Screening
The second stage of the semi-classical analysis involves examining the propagation and interactions of the quantum fluctuations $\delta\phi(x)$ on top of the background $\overline{\phi}(x)$ in the regime where $\mathcal{E}\gg\Lambda_{\ast}$. It is essential to verify that these quantum fluctuations do not spoil the semi-classical approximation, analogous to the treatment of classical potentials in inflationary models [162, 166, 167]. We will discuss that quantum corrections are naturally suppressed by the quantum version of Vainshtein screening.
For convenience, we decompose the field as $\phi(x)=\overline{\phi}(x)+\delta\phi(x)$, and similarly decompose the kinetic variable $X=\overline{X}+\delta X$, where
$\overline{X}\equiv\frac{\partial^{\mu}\overline{\phi}\partial_{\mu}\overline{\phi}}{2\Lambda_{\ast}^{4}},\quad\text{and}\quad\delta X\equiv\frac{\partial^{\mu}\delta\phi\,\partial_{\mu}\delta\phi+2\partial^{\mu}\overline{\phi}\partial_{\mu}\delta\phi}{2\Lambda_{\ast}^{4}}.$ (2.18)
The Lagrangian density $\mathcal{L}_{X}$ in Eq. (2.7) can then be expressed in terms of the fluctuations $\delta\phi(x)$. Expanding the kinetic function in a Taylor series yields
$\mathcal{K}(\overline{X}+\delta X)=\mathcal{K}(\overline{X})+\mathcal{K}^{(1)}(\overline{X})\,\delta X+\frac{\mathcal{K}^{(2)}(\overline{X})}{2}\,\delta X^{2}+\mathcal{O}\left(\delta X^{3}\right),$ (2.19)
where $\mathcal{K}^{(n)}(X)\equiv d^{n}\mathcal{K}(X)/dX^{n}$. Given our specific choice of $\mathcal{K}(X)$ in Eq. (2.8), this second-order expansion is exact, since
$\mathcal{K}^{(1)}(X)=1+2c_{2}X,\quad\mathcal{K}^{(2)}(X)=2c_{2},\quad\text{and}\quad\forall n>2,\,\mathcal{K}^{(n)}(X)=0.$ (2.20)
Kinetic Term:
From the original Lagrangian density $\mathcal{L}_{X}$ in Eq. (2.7), we derive the kinetic term for the massless fluctuations $\delta\phi(x)$, which takes the form:
$\mathcal{L}_{\mathrm{kin}}=\frac{\mathcal{Z}_{\phi}^{\mu\nu}}{2}\,\partial_{\mu}\delta\phi\partial_{\nu}\delta\phi,\quad\text{where}\quad\mathcal{Z}_{\phi}^{\mu\nu}\equiv\mathcal{K}^{(1)}(\overline{X})\,\eta^{\mu\nu}+\frac{\mathcal{K}^{(2)}(\overline{X})}{\Lambda_{\ast}^{4}}\,\partial^{\mu}\overline{\phi}\partial^{\nu}\overline{\phi},$ (2.21)
with $\eta_{\mu\nu}$ denoting the Minkowski metric. Owing to the spherical symmetry of the background $\overline{\phi}(x)\equiv\overline{\phi}(r)]$, the kinetic term can be recast as
$\mathcal{L}_{\mathrm{kin}}=\frac{Z_{\phi}(x)}{2}\left[(\partial_{t}\delta\phi)^{2}-(\partial_{\Omega}\delta\phi)^{2}-B(x)\cdot(\partial_{r}\delta\phi)^{2}\right],$ (2.22)
where the angular derivative term is given by
$(\partial_{\Omega}\delta\phi)^{2}\equiv\left(\frac{\partial_{\theta}\delta\phi}{r}\right)^{2}+\left(\frac{\partial_{\varphi}\delta\phi}{r\sin\theta}\right)^{2},$ (2.23)
and the introduced non-vanishing functions are defined as
$Z_{\phi}(x)\equiv\mathcal{K}^{(1)}(\overline{X}),\quad B(x)=\dfrac{\mathcal{K}^{(1)}(\overline{X})+2\overline{X}\,\mathcal{K}^{(2)}(\overline{X})}{\mathcal{K}^{(1)}(\overline{X})},$ (2.24)
with $Z_{\phi}>0$ required to avoid ghost instabilities [3]. It is important to note that the kinetic term is generally manifestly anisotropic. This anisotropy is acceptable, as the presence of the external source $J(x)$ inherently breaks the manifest Lorentz invariance of the original Lagrangian density $\mathcal{L}_{X}$. Below, we discuss the distinct regimes of the background solution in Eq. (2.16), based on our choice of kinetic function in Eq. (2.8).
In the Vainshtein core ($r\ll R_{V}$), we find $B(r)\to 3$, and the ghost-free condition imposes
$Z_{\phi}(r)\underset{|\overline{\phi}^{\prime}|\gg\Lambda_{\ast}^{2}}{\sim}-c_{2}\left[\frac{\overline{\phi}^{\,\prime}(r)}{\Lambda_{\ast}^{2}}\right]^{2}>0\quad\Longrightarrow\quad c_{2}=-1,$ (2.25)
thus recovering the same condition on $c_{2}$ required for the existence of $\overline{\phi}(r\ll R_{V})$. Failure to satisfy this condition would indicate an internal inconsistency within this k-essence model. Zooming in on a specific point $\mathbf{r_{0}}\equiv(r_{0},\theta_{0},\varphi_{0})$ within the Vainshtein core, with a resolution of $\Delta\mathbf{r}\equiv(\Delta r,\Delta\theta,\Delta\varphi)$, we obtain
$\sqrt{\frac{1}{Z_{\phi}(r_{0}+\Delta r)}}\underset{\Delta r\,\ll\,r_{0}}{=}\sqrt{\frac{1}{Z_{\phi}(r_{0})}}+\Delta r\cdot\left.\frac{d}{dr}\left[\sqrt{\frac{1}{Z_{\phi}(r)}}\right]\right|_{r=r_{0}}+\mathcal{O}\left(\frac{\Delta r^{2}}{r_{0}^{2}}\right),$ (2.26)
where
$\Delta r\cdot\left.\frac{d}{dr}\left[\sqrt{\frac{1}{Z_{\phi}(r)}}\right]\right|_{r=r_{0}}\sim\left(\frac{r_{0}}{R_{V}}\right)^{\frac{2}{3}}\frac{\Delta r}{r_{0}}\ll 1,$ (2.27)
such that only the leading-order term is retained. Consequently, we can perform the field-strength renormalization of the fluctuation field to absorb the isotropic factor $Z_{\phi}(r_{0})$ in the kinetic term, leading to a Taylor expansion that yields
$\mathcal{L}_{\mathrm{kin}}\sim\frac{1}{2}\left[(\partial_{t}\delta\phi_{Z})^{2}-(\partial_{\Omega}\delta\phi_{Z})^{2}-3(\partial_{r}\delta\phi_{Z})^{2}\right],\quad\text{with}\quad\delta\phi_{Z}(x)\equiv\sqrt{Z_{\phi}(r_{0})}\,\delta\phi(x).$ (2.28)
This field-strength renormalization is essential for determining the regime of validity of the EFT for the fluctuations, as we will discuss when analyzing the interaction terms.
In the linear ($r\gg R_{V}$) and transition ($r\sim R_{V}$) regimes, we have $Z_{\phi}(r)\sim B(r)\sim 1$, and the ghost-free condition remains consistent with $c_{2}=-1$. While it is possible to introduce the renormalized field $\delta\phi_{Z}(x)\equiv\sqrt{Z_{\phi}}\,\delta\phi(x)$, this field-strength renormalization does not significantly modify the interaction scale, as we will demonstrate.
Interaction Terms:
The cubic and quartic interaction terms for $\delta\phi(x)$ are given by
$\mathcal{L}_{\mathrm{int}}=-\frac{\partial^{\nu}\overline{\phi}\partial_{\nu}\delta\phi\,\partial^{\mu}\delta\phi\,\partial_{\mu}\delta\phi}{\Lambda_{\ast}^{4}}-\frac{\left(\partial^{\mu}\delta\phi\,\partial_{\mu}\delta\phi\right)^{2}}{4\Lambda_{\ast}^{4}},$ (2.29)
respectively. It is important to emphasize that the shift symmetry remains preserved by the Lagrangian density of the fluctuations $(\delta\phi\mapsto\delta\phi+\phi_{c})$. We now focus on the distinct regimes of the background solution, as outlined in Eq. (2.16), using our chosen kinetic function in Eq. (2.8).
In the non-linear regime ($r\ll R_{V}$), the interaction terms exhibit the following asymptotic forms:
$\mathcal{L}_{\mathrm{int}}\sim\dfrac{\partial_{r}\delta\phi_{Z}\,\partial^{\mu}\delta\phi_{Z}\,\partial_{\mu}\delta\phi_{Z}}{\overline{\Lambda_{\ast}}^{2}}-\dfrac{\left(\partial^{\mu}\delta\phi_{Z}\,\partial_{\mu}\delta\phi_{Z}\right)^{2}}{4\overline{\Lambda_{\ast}}^{4}},$ (2.30)
where, after field-strength renormalization, the effective interaction scale at which the coupling becomes of $\mathcal{O}(1)$ is
$\overline{\Lambda_{\ast}}(r_{0})\equiv\sqrt{Z_{\phi}(r_{0})}\,\Lambda_{\ast},$ (2.31)
which depends explicitly on $r_{0}$. Based on the background solution in the Vainshtein core, as provided in Eq. (2.16), and the expression for the renormalization factor $Z_{\phi}(r_{0})$ in Eq. (2.25), we can draw the following conclusions:
  • $\spadesuit$
    Deep within the Vainshtein core, $\overline{\Lambda_{\ast}}(r_{0})\gg\Lambda_{\ast}$. This indicates that the perturbative cutoff $\overline{\Lambda_{\ast}}$ is blueshifted relative to the original (naive) cutoff $\Lambda_{\ast}$ in $\mathcal{L}_{X}$. Consequently, the regime of validity of the EFT is ‘redressed’ by Vainshtein screening: the interaction scale for fluctuations about the classicalon background becomes $\overline{\Lambda_{\ast}}(r_{0})\gg\Lambda_{\ast}$, rather than $\Lambda_{\ast}$. The stronger the blueshift effect, the more suppressed the self-interactions in Eq. (2.30). This suppression represents another manifestation of Vainshtein screening: self-interactions among quantum fluctuations are weakened. As a corollary, loop corrections are also suppressed, as demonstrated for k-essences in Refs. [168, 153, 151] and for Galileons in Refs. [143, 144]. In essence, quantum effects do not disrupt the classical background in the Vainshtein core due to the screening mechanism.
  • $\spadesuit$
    For a fixed $\mathbf{r_{0}}$, the factor $Z_{\phi}(r_{0})$ increases with $\mathcal{E}$, enhancing the blueshift effect for larger values of $\mathcal{E}$. It is worth noting that the volume of the Vainshtein core also scales as $R_{V}^{3}$.
  • $\spadesuit$
    For a given value of $\mathcal{E}$, the factor $Z_{\phi}(r_{0})$ increases as $r_{0}\to 0$, strengthening the blueshift effect as one approaches the source. Notably, $\overline{\phi}^{\,\prime}(r)$ diverges for $r_{0}\to 0$, reflecting the pointlike nature of the external source $J(x)$.
Applying the same methodology, it is evident that in the linear ($r\gg R_{V}$) and transition ($r\sim R_{V}$) regimes, the interaction scale for fluctuations about the classicalon background remains $\Lambda_{\ast}$, since the fluctuation fields undergo negligible field-strength renormalization ($Z_{\phi}\sim 1$). Fortunately, quantum fluctuations are expected to remain under EFT control in this case, which is why the blueshift effect is critical only within the Vainshtein core.
In summary, Vainshtein screening corresponds to the dominance of the kinetic self-interaction over the kinetic term of the k-essence field. The Vainshtein core thus defines a spatial region in which $\phi$-bosons cannot propagate as weakly interacting particles, as non-linear interactions prevail [137, 151]. From the perspective of background fluctuations $\delta\phi(x)$, their kinetic term acquires a large renormalization due to the background for $r\ll R_{V}$; the Vainshtein mechanism is therefore a form of screening by inertia [3]. After field strength renormalization, this corresponds to fluctuations $\delta\phi_{Z}(x)$ that behave effectively as quasi-free particles in the limit $r_{0}\to 0$ via the blueshift of the interaction scale $\Lambda_{\ast}$.
2.2.4 Classicalization from Kinetic Self-interactions
The foregoing analysis demonstrates that classicalon solutions arise in response to a sharply localized external energy-momentum source. This ensures that any attempt to localize quanta coupled to $\phi(x)$ within $R_{V}$ will generate a Vainshtein core surrounding the source, thereby preventing the probing of shorter length scales through the emission of a hard k-essence boson.
In a toy model with only k-essence bosons, there is no external source: the localized energy-momentum source corresponds to the momentum exchange of scattered $\phi$-bosons themselves, with a center-of-mass energy $\sqrt{s}\gg\Lambda_{\ast}$, where the kinetic self-interaction acts as a self-sourcing term that prevents the probing of length scales below the classicalization radius $R_{\circledast}$. We now examine how the previous discussion is altered.
Self-sourcing:
As discussed in Ref. [108], the radius $R_{\circledast}$ can be obtained by examining the static scenario in which a field configuration $\phi_{0}(r)$, characterized by an energy $M_{\circledast}\sim\sqrt{s}$, is confined within a sphere of radius $R_{\circledast}$, the latter representing the typical scale of variation for $\phi_{0}(r)$. A dimensional analysis yields
$M_{\circledast}\sim\int_{r<R_{\circledast}}d^{3}\mathbf{r}\left(\partial\phi_{0}\right)^{2}\sim\left.\Omega r\phi_{0}^{2}\right|_{r=R_{\circledast}}.$ (2.32)
The localized field configuration $\phi_{0}(r)$ of size $R_{\circledast}$ then serves as a source for the field $\phi(x)$ itself, with an integrated value parametrically given by
$\left|\frac{\mathcal{E}_{0}}{\Lambda_{\ast}}\right|\sim\left|\int_{r<R_{\circledast}}d^{3}\mathbf{r}\ \frac{\delta}{\delta\phi_{0}}\left[\frac{\left(\partial\phi_{0}\right)^{4}}{\Lambda_{\ast}^{4}}\right]\right|\sim\left.\left|\frac{\Omega}{r}\cdot\frac{\phi_{0}^{3}}{\Lambda_{\ast}^{4}}\right|\right|_{r=R_{\circledast}}.$ (2.33)
An analysis analogous to that presented in Section 2.2.2, in which the pointlike source is replaced by this extended source, yields
$\forall\,r>R_{\circledast}\,,\quad\phi(r)\sim-\frac{1}{\Omega r}\cdot\frac{\mathcal{E}_{0}}{\Lambda_{\ast}},$ (2.34)
which exhibits a Coulomb-like tail outside the source. This leads to the classicalization radius $R_{\circledast}$ for the k-essence field $\phi(x)$, where the kinetic self-interaction begins to dominate the kinetic term:
$R_{\circledast}\equiv\ell_{\ast}\left(\frac{M_{\circledast}}{\Omega\Lambda_{\ast}}\right)^{\frac{1}{3}}\gg\ell_{\ast},\quad\text{with}\quad M_{\circledast}\gg\Lambda_{\ast}.$ (2.35)
This result has been explicitly confirmed through the study of spherical wavepacket collapse, both analytically [109, 67] and numerically [113, 115]. These studies provide evidence that the wavepacket propagates freely until reaching the distance $R_{\circledast}$, at which point scattering effects prevent the resolution of smaller length scales.
From this discussion, it becomes clear how k-essence achieves self-UV-completion through classicalization during a hard scattering process with $\sqrt{s}\gg\Lambda_{\ast}$: a classicalon of radius $R_{\circledast}$ and mass $M_{\circledast}\sim\sqrt{s}$ is formed. The greater the energy injected by the scattering process, the larger the classicalon and the stronger the blueshift effect within the Vainshtein core for the UV modes. The behavior of $R_{\circledast}$ is analogous to that of the Schwarzschild radius for a BH of mass $M_{\circledast}\sim\sqrt{s}$,
$R_{S}\equiv\ell_{P}\left(\dfrac{M_{\circledast}}{\Omega\Lambda_{P}}\right)\gg\ell_{P},\quad\text{where}\quad M_{\circledast}\gg\Lambda_{P},$ (2.36)
beyond which gravitational non-linearities become significant. In Fig. 5, we compare the efficiency of the classicalization phenomenon between gravity and k-essence for $\ell_{\ast}=\ell_{P}$. As discussed in Ref. [67], gravity represents the most efficient classicalizer, as the Schwarzschild radius in Eq. (2.36) grows more rapidly than $R_{\circledast}$ with increasing $M_{\circledast}$.
Figure 5: A comparison between gravity and massless k-essence with respect to the growth rate of the classicalon radius $R_{\circledast}$ as a function of its mass $M_{\circledast}$. To ensure a meaningful comparison, the classicalization length $\ell_{\ast}\equiv 1/\Lambda_{\ast}$ is fixed to the Planck length $\ell_{P}$ in both scenarios. The plot indicates that $R_{\circledast}$ increases more rapidly with increasing $M_{\circledast}$ in gravity than in k-essence.
Dimensional Analysis:
It is illuminating to demonstrate that the parametric form of $R_{\circledast}$ can be derived through dimensional analysis alone [108]. To this end, we adopt a system of units [169] that restores the $\hbar$-dependence while retaining $c=1$, thereby distinguishing units of energy $(E)$ and length $(L)$. Our focus is on an inclusive scattering process involving 2 $\phi$-bosons with a center-of-mass energy $\sqrt{s}$ and an impact parameter $b\sim 1/\sqrt{s}$. The action for the field $\phi(x)$ is expressed as
$S_{X}=\int d^{4}x\ \mathcal{L}_{X},\quad\mathcal{L}_{X}=\frac{1}{2}\,\partial^{\mu}\phi\partial_{\mu}\phi+\frac{G_{\phi}}{2}\left(\partial^{\mu}\phi\partial_{\mu}\phi\right)^{2}.$ (2.37)
The dimensionalities of the relevant quantities are as follows:
$[\partial]=L^{-1},\quad[d^{4}x]=L^{4},\quad[S_{X}]=[\hbar]=EL,\quad[\mathcal{L}_{X}]=EL^{-3},$
$[\phi]=E^{\frac{1}{2}}L^{-\frac{1}{2}},\quad[G_{\phi}]=E^{-1}L^{3},\quad[\sqrt{s}]=E,$ (2.38)
where the dimensions in natural units $(\hbar=1)$ are recovered by setting $E=L^{-1}$.
Relationships among physical quantities should depend solely on the parameters $\hbar$, $G_{\phi}$, and $\sqrt{s}$. From these, one can construct: $(i)$ a unique quantum length associated exclusively with the system’s kinematics,
[…]
which corresponds to the de Broglie wavelength of the source (i.e., the incoming particles); $(ii)$ a unique quantum length associated solely with the kinetic self-coupling,
$\ell_{\ast}\equiv\left|\hbar G_{\phi}\right|^{\frac{1}{4}},$ (2.40)
which defines the length scale at which this interaction becomes strongly coupled; and $(iii)$ a unique classical length,
$r_{c}\equiv\left|\sqrt{s}G_{\phi}\right|^{\frac{1}{3}},$ (2.41)
[…]
Quantum Criticality:
As a semi-classical state, a classicalon can be characterized as a coherent state—a quantum state with a large occupation number $N_{\circledast}\gg 1$ of k-essence bosons. By definition,
$N_{\circledast}\equiv\frac{M_{\circledast}}{\omega}\sim\left(\frac{M_{\circledast}}{\Lambda_{\ast}}\right)^{\frac{4}{3}},$ (2.42)
where $\omega$ is the average energy of a constituent boson, see Eqs. (2.6) and (2.35). These soft bosons interact through a quantum coupling $\alpha$, which can be derived from the Lagrangian density $\mathcal{L}_{X}$ in Eq. (2.9):
$\alpha\equiv\left(\frac{\omega}{\Lambda_{\ast}}\right)^{4}\sim\left(\frac{M_{\circledast}}{\Lambda_{\ast}}\right)^{-\frac{4}{3}}.$ (2.43)
One can then introduce a collective coupling for the classicalon state,
$\alpha_{c}\equiv\alpha N_{\circledast}\sim 1,$ (2.44)
a hallmark of a saturon [158, 80]. This result reflects the quantum criticality of the $N_{\circledast}$-boson system, marking a quantum critical point at which collective interactions among the classicalon constituents become significant. Despite the weak coupling between individual bosons ($\alpha\ll 1$), this critical point signals the onset of non-perturbative collective effects [121, 124]. Therefore, classicalization can be interpreted as the attainment of a fixed point for the collective coupling [^15] This holds provided that a mass term does not prevent the classicalization radius from exceeding the corresponding Compton wavelength (see Section 3.2.1 for an example).—exhibiting critical behavior—rather than for the individual couplings between the constituents.
2.3 New Landscape Islands in Terra Incognita
Self-UV-completion by classicalization opens the possibility of a new Landscape of QFTs, potentially building a bridge to quantum gravity. However, many open issues certainly remain. In particular, this topic does not possess the same degree of maturity as textbook QFT [1, 17], and the reader may have several concerns about the viability of such a proposal. In the following, we review several criticisms that have been raised in the community, and their proposed resolutions, without claiming any originality. The objective is to briefly summarize and coherently present the relevant literature, which is often dispersed across various topics and communities.
2.3.1 Trouble with Asymptotic States? Classicalons vs Hadrons
The concept of asymptotic darkness—that is, the dominance of BH states in ultra-Planckian hard scatterings of particles—has occasionally been criticized on the grounds of the definition of asymptotic states. Since classicalization generalizes this phenomenon to interactions beyond gravity, such criticisms are equally relevant. For example, in Section 2 of Ref. [170], it is argued that the notion of gravitons as asymptotic states in the ultra-Planckian regime may be ill-defined due to non-perturbative graviton self-interactions [^16] This criticism is unrelated to the legitimate discussions concerning IR divergences and the soft dressing of asymptotic states. Such subtleties arise even in the absence of classicalization, and we have no further insights to contribute on this matter., analogous to how quarks and gluons cannot be defined as asymptotic states in QCD because they are confined within hadrons.
Despite the fact that both BH production in GR and hadronization in QCD arise from non-perturbative dynamics, the comparison between these 2 phenomena is not appropriate for critiquing classicalization. To clarify this distinction, let us consider the scattering of 2 massive particles, each with mass $M\ll\Lambda_{P}$, in the context of GR, rather than massless particles. For an ultra-Planckian hard scattering characterized by $\sqrt{s}\gg\Lambda_{P}$, one expects the production of a BH with mass $M_{\circledast}\sim\sqrt{s}$. A naive question arises: Why are the initial particle states well-defined as particles (rather than BHs) if they themselves possess ultra-Planckian energies $E\gg\Lambda_{P}$, similar to how quarks cannot be defined as asymptotic states in QCD? The key point [171] is that energy and momentum are not Lorentz-invariant quantities; one can perform a Lorentz boost to study the particle in its rest frame, where $E\ll\Lambda_{P}$. The Lorentz-invariant quantity for an asymptotic particle is its mass $M\ll\Lambda_{P}$. While such a particle can emit soft gravitons in its rest frame, this emission occurs within the weak-field regime, where BH formation does not take place.
In the case of scattering between 2 particles of mass $M\ll\Lambda_{P}$, 2 Lorentz-invariant and independent kinematic quantities can be defined: the Mandelstam variables [^17] The Mandelstam variable tt—for a general process involving 2 incoming particles—is defined in the conventional manner for a would-be $2\to 2$ scattering process. $s$ and $t$. If $\sqrt{s}\gg\Lambda_{P}$ but $\sqrt{-t}\ll\Lambda_{P}$, with the impact parameter $b\gg R_{S}\sim\ell_{P}^{2}\sqrt{s}$, the system does not form a BH; instead, it can be analyzed through the eikonal resummation of soft graviton exchange [24, 25]. BH formation is expected only when $\sqrt{s}\gg\Lambda_{P}$ and $b\ll R_{S}$.
In contrast, for light quarks in QCD, hadronization results from strong coupling in the IR regime: in its rest frame, an ‘asymptotic’ quark is strongly coupled to the gluon field and will inevitably form a hadron. Therefore, although non-perturbative effects are fundamental in both GR and QCD, their underlying mechanisms are fundamentally distinct. This discussion can be extended to massless particles as well (which lack a rest frame), with the critical observation being that energy and momentum are not Lorentz-invariant quantities.
2.3.2 Beyond Spherical Symmetry: Classicalization vs Eikonal Regimes
In the previous sections, our analysis has focused exclusively on classicalization involving perfectly spherical sources. However, Ref. [172] demonstrates that classicalon formation fails to occur in scattering configurations that deviate substantially from spherical symmetry, thereby calling into question the viability of UV-completing an EFT via this mechanism. The efficiency of classicalization is maximized for a perfectly spherical source, a feature associated with the diminished effectiveness of Vainshtein screening in non-spherical configurations [173, 174].
In fact, classicalization is the phenomenon responsible for unitarizing a scattering process with $\sqrt{s}\gg\Lambda_{\ast}$ and an impact parameter $b\lesssim R_{\circledast}$ [67]. This scenario represents the configuration closest to spherical symmetry. A significant deviation from spherical symmetry, characterized by $b\gg R_{\circledast}$, implies that unitarization is instead expected to occur through the eikonal resummation of the large number of soft mediators exchanged in the $t$-channel between the scattered particles. This mechanism is well documented in the literature on the gravitational $S$-matrix [24, 25] and is distinct from Vainshtein screening, which can also be investigated via the resummation of Feynman diagrams [175]. Consequently, depending on the hierarchy between $b$ and $R_{\circledast}$, the non-perturbative mechanism underlying unitarity should exhibit a smooth transition between the eikonal and classicalization regimes [67].
2.3.3 Causality and Positivity Bounds: The Fate of the Fate
As discussed in Ref. [147], certain non-gravitational EFTs with apparently local and Lorentz-invariant actions can give rise to superluminal propagation on specific semi-classical backgrounds, even within their regime of validity. It is important to emphasize that superluminality, in itself, does not constitute a pathology. However, it has been argued that, within such EFTs, it is possible to construct closed timelike curves (CTCs) at the classical level, thereby challenging the conventional notion of causality. The underlying reasoning is that a UV-completion consistent with the standard axioms of QFT—including Lorentz invariance, unitarity, analyticity, and locality—imposes positivity bounds on certain coefficients of the EFT operators to prevent such issues [147]. These bounds can also be derived from dispersion relations [163, 147, 164, 165]. It is well-established, however, that the k-essence theory with $c_{2}=-1$ [163, 147, 116, 164, 165] and the Galileons theories [176, 177, 178, 179] violate these positivity bounds, thereby raising doubts about the validity of classicalization. The issue of superluminality and CTCs for k-essence and Galileon fields has been highlighted in Refs. [180, 181, 182, 183, 184, 185].
The critique based on positivity bounds can be challenged by examining its 2 foundational pillars:
$(i)$ Superluminality:
Drawing an analogy with chronology protection arguments in GR [186, 187], several studies for both k-essence and Galileon theories [188, 189, 190] have argued that strong quantum backreaction renders CTCs sensitive to UV physics, thereby indicating that such a pathological background lies beyond the regime of control within the EFT. This observation aligns with expectations for theories that achieve self-UV-completion through classicalization, as discussed in Ref. [116]. Qualitatively, classicalization introduces the concept of a fundamental limit on spacetime resolution associated with a given field. As a result, light cones appear fuzzy, and the notion of a pointlike event loses its meaning. It is therefore unsurprising that the traditional concept of microcausality must be reconsidered. Further exploration of the distinction between micro- and macro-causality in theories exhibiting Vainshtein screening can be found in Refs. [133, 134, 191]. The conclusion, based on these studies, is that k-essence and Galileon theories do not result in dramatic violations of causality, such as the possible existence of CTCs.
$(ii)$ Dispersion Relations:
Classicalization embodies the notion of a minimal length scale that an interaction can probe and is thus non-perturbatively non-local, despite the apparent locality of the k-essence Lagrangian density in Eq. (2.9). This can be formalized within the axiomatic framework of QFT using Jaffe’s classification of strictly localizable, quasi-local, and non-localizable fields, specifically by examining the ultraviolet behavior of the spectral density functions [192, 193, 194, 134, 195]. The Wightman formulation, causality, and the standard properties of the $S$-matrix for theories with non-localizable fields have been studied in Refs. [196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210]. Positivity bounds are typically derived from dispersion relations for textbook QFTs, which deal with tempered localizable fields, a subclass of strictly localizable fields [195], whereas classicalizing QFTs involve non-localizable fields [134, 135]. The status of positivity bounds on QFTs with non-localizable fields is discussed in Refs. [211, 195, 135, 212], and the results for classicalizing theories are consistent with the violation of the bounds obtained by assuming tempered localizable fields [195, 135], e.g., $c_{2}=-1$. The properties of non-localizable fields, whose spectral density functions grow faster than those of strictly localizable fields, explain why classicalization appears at odds with any interpretation in a Wilsonian perspective [148]. We also stress that this is the reason the classicalization proposal for scalars is fundamentally different from the controversial Higgsplosion scenario [213, 214, 215, 216, 217] in the framework of the standard electroweak theory, involving only strictly localizable fields [218, 219], while both deal with multi-scalar boson production at high energies.
The conclusion of this discussion is that the non-local features of classicalization invalidate the usual mathematical hypotheses used to derive positivity bounds from dispersion relations in local QFTs, while exhibiting a self-protection mechanism against the ability to observe violations of causality.
2.3.4 Non-renormalizable Operators Demystified
In the models under consideration, classicalization emerges through operators—referred to as UV-screeners—that belong to the infinite class of non-renormalizable operators within the conventional EFT framework. Since this phenomenon becomes apparent in hard scattering processes with $\sqrt{s}\gg\Lambda_{\circledast}$, practitioners of EFTs may find the interpretation of these non-renormalizable operators within the classicalization paradigm unclear. To clarify this point, it is worthwhile to first review the UV origin of non-renormalizable operators in a standard EFT [5].
Ultraviolet Remnants:
The action of an EFT can be expressed as
$S_{\mathrm{EFT}}=\int d^{4}x\left(\mathcal{L}_{\mathrm{R}}+\mathcal{L}_{\mathrm{NR}}\right),$ (2.45)
where $\mathcal{L}_{\mathrm{R}}$ and $\mathcal{L}_{\mathrm{NR}}$ denote the renormalizable and non-renormalizable parts of the Lagrangian density, respectively. Within the Wilsonian EFT framework, the latter are termed ‘irrelevant’ because, within the regime of validity of the EFT, they are suppressed by powers of a heavy interaction scale $\Lambda$ according to
$\mathcal{L}_{\mathrm{NR}}=\sum_{k=1}^{+\infty}\dfrac{c_{k}}{\Lambda^{d_{k}-4}}\,\mathcal{O}^{(k,\,d_{k})}_{\mathrm{NR}}$, (2.46)
where $\mathcal{O}^{(k,d_{k})}_{\mathrm{NR}}$ represents an operator of dimension $d_{k}>4$, and the coefficients $c_{k}$ are naturally of $\mathcal{O}(1)$ unless suppressed by selection rules from (approximate) symmetries. In the Wilsonian approach, this infinite tower of operators arises from integrating out heavy fields.
To illustrate, consider a toy model featuring 1 real scalar field $\Phi(x)$ of mass $M_{\Phi}$ and 1 massless Dirac fermion $\psi(x)$. The Lagrangian density, including a Yukawa coupling $y_{\psi}$, is given by
$\mathcal{L}_{\mathrm{UV}}[\Phi,\bar{\psi},\psi]=i\bar{\psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\psi-\dfrac{1}{2}\,\Phi(\square+M_{\Phi}^{2})\Phi+y_{\psi}\,\bar{\psi}\psi\Phi,$ (2.47)
with the Dirac matrices $\gamma^{\mu}$, and the derivative operator defined as $\overleftrightarrow{\partial_{\mu}}=\overrightarrow{\partial_{\mu}}-\overleftarrow{\partial_{\mu}}$. An EFT involving only the light field $\psi(x)$ can be obtained by integrating out the heavy field $\Phi(x)$. The effective action $S_{\mathrm{EFT}}$ is formally non-perturbatively defined as
$e^{iS_{\mathrm{EFT}}}=\int[\mathcal{D}\Phi]\ e^{iS_{\mathrm{UV}}[\Phi,\,\bar{\psi},\,\psi]},$ (2.48)
where $S_{\mathrm{UV}}$ is the action associated with $\mathcal{L}_{\mathrm{UV}}$. At tree level, $\Phi$ can be integrated out using its Euler-Lagrange equation, which amounts to the substitution
$\Phi\mapsto y_{\psi}\left(\dfrac{1}{\square+M_{\Phi}^{2}}\right)\bar{\psi}\psi,$ (2.49)
yielding a non-local effective Lagrangian density:
$\mathcal{L}_{\mathrm{EFT}}^{\mathrm{Nloc}}[\bar{\psi},\psi]=i\bar{\psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\psi+\frac{y_{\psi}^{2}}{2}\,\bar{\psi}\psi\left(\dfrac{1}{\square+M_{\Phi}^{2}}\right)\bar{\psi}\psi.$ (2.50)
This EFT is equivalent to the original theory only at tree level [^18] More generally, a systematic method for integrating out fields in perturbation theory is the background field method [220]. for the $\Phi$-field, as the non-local form factor in the quartic term encodes the tree-level propagation of the massive $\Phi$-particle. For Euclidean momenta $p_{E}^{2}\ll M^{2}$, a Taylor expansion of this form factor—known as the operator product expansion (OPE)—produces a local EFT Lagrangian density:
$\mathcal{L}_{\mathrm{EFT}}^{\mathrm{loc}}[\bar{\psi},\psi]=i\bar{\psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\psi+\frac{y_{\psi}^{2}}{2}\left[\frac{\bar{\psi}\psi\bar{\psi}\psi}{M_{\Phi}^{2}}-\frac{\bar{\psi}\psi\square\bar{\psi}\psi}{M_{\Phi}^{4}}+\mathcal{O}\left(\frac{\square^{2}}{M_{\Phi}^{6}}\right)\right],$ (2.51)
where the tower of effective operators is truncated at the appropriate order in $\square/M^{2}$ via power counting. These non-renormalizable operators [^19] While this terminology is standard, it is not accurate, as the EFT remains perturbatively renormalizable at any given order in perturbation theory. thus represent the low-energy remnants of the $\Phi$-propagator, with the interaction scale $\Lambda$ set by the heavy mass $M$ if $y_{\psi}\sim 1$.
The role of non-renormalizable operators— referred to as ‘UV-remnants’—is evident from this example: they provide an approximate description of the quantum fluctuations of heavy particles and are not fundamental. Crucially, these operators are treated perturbatively and must remain so, even when seeking semi-classical solutions within an EFT (which are non-perturbative in couplings but perturbative in the EFT expansion). This perturbative treatment ensures that EFTs remain free of ghost instabilities, in contrast to generic higher-derivative theories [221, 166, 222, 223].
Ultraviolet Sreeners:
In the context of classicalization, the non-renormalizable operators (UV-screeners) responsible for this phenomenon must be treated non-perturbatively. To maintain consistency with the discussion about UV-remnants, this implies that their fundamental origin cannot be attributed to the OPE of form factors coming from integrating out UV degrees of freedom [110, 116, 103]. To clarify this distinction, the non-renormalizable operators in the EFT can be decomposed as follows:
$\mathcal{L}_{\mathrm{NR}}=\sum_{i\geq 1}\dfrac{\kappa_{i}}{\Lambda_{\ast}^{d_{i}-4}}\,\mathcal{O}^{(i,\,d_{i})}_{\mathrm{scr}}+\sum_{j=1}^{+\infty}\dfrac{c_{j}}{M^{d_{j}-4}}\,\mathcal{O}^{(j,\,d_{j})}_{\mathrm{rem}}$, (2.52)
where the operators $\mathcal{O}^{(i,d_{i})}_{\mathrm{scr}}$ and $\mathcal{O}^{(j,d_{j})}_{\mathrm{rem}}$ represent UV-screeners and UV-remnants of dimension $d_{i}>4$ and $d_{j}>4$, respectively. Here, $\kappa_{i}$ and $c_{j}$ are dimensionless couplings, $\Lambda_{\ast}$ denotes the classicalization scale, and $M$ is the mass scale associated with the next layer of UV degrees of freedom that have been integrated out. Given that classicalization is intended to UV-complete the theory within the energy range between $\Lambda_{\ast}$ and $M$, the hierarchy $\Lambda_{\ast}\ll M$ must hold. Note that the number of UV-screeners may be infinite if they are defined through arbitrary functions of the fields and their derivatives, like $\mathcal{K}(X)$ in Eq. (2.7).
For energies $E\ll\Lambda_{\ast}$, the non-perturbative dynamics of UV-screeners—including fuzzyons and the infinite tower of classicalons—can be integrated out. This procedure yields a non-local EFT Lagrangian density, which can then be expanded into a tower of UV-remnants via the OPE. Thus, below the scale $\Lambda_{\ast}$, an EFT can be constructed using conventional methods, despite an apparent violation of traditional positivity bounds.
Since UV-screeners govern the classicalization dynamics, they must be treated non-perturbatively, and their selection must ensure the absence of physical ghost-like degrees of freedom in the spectrum. This challenge is well documented in the modified gravity literature [3], where k-essence and Galileon theories are specifically constructed to avoid such pathologies [^20] For a broader discussion of these constructions (in the context of dark energy theories), see the review [224].. This requirement represents a fundamental distinction from UV-remnants.
As a final remark on the interpretation of UV-screeners within the EFT framework, the reader may question how the values of the couplings $\kappa_{i}$ align with the renormalization program. A detailed analysis of the perturbative renormalization of the EFT governing the fluctuations around the classicalon background lies beyond the scope of this article (see Ref. [153] for an attempt in this direction). Here, we offer only preliminary considerations on the subject. Within the classicalon radius, the running of these couplings is effectively frozen, as the dynamics in this region is predominantly classical, as demonstrated in Ref. [151]. Since quantum corrections are suppressed, no large logarithms should require resummation. Consequently, one should be able to choose the renormalized values of the couplings $\kappa_{i}$ to match those of the classical Lagrangian $(\hbar\to 0)$, with, e.g., a subtraction point specified by the background amplitude $\overline{\phi_{0}}^{\prime}\equiv\Lambda_{\ast}^{2}$. The counterterms should then be employed to absorb the UV divergences, as usual in the background field method [1]. This also implies that if an operator vanishes classically, it is incorporated with a renormalized coupling that is zero in the Lagrangian density, yet accompanied by a non-vanishing counterterm to address the UV divergences [153].
3 From K-essence to K-chameleon
Let us revisit the massless k-essence model analyzed in Section 2.2. For physical applications, scalar fields are typically introduced with a potential $V(\phi)$, which explicitly breaks the shift symmetry. In the presence of a localized source, the Euler-Lagrange equation (2.12) takes the modified form
$\square\phi+\frac{c_{2}}{\Lambda_{\ast}^{4}}\,\partial^{\mu}\left(\partial^{\nu}\phi\,\partial_{\nu}\phi\,\partial_{\mu}\phi\right)+\frac{dV}{d\phi}=-\dfrac{\mathcal{E}}{\Lambda_{\ast}}\,\delta^{(3)}(\mathbf{r}),$ (3.1)
where $c_{2}=-1$ and the left-hand side is no longer a total divergence. Our objective is to study how this potential affects Vainshtein screening—and, consequently, the conditions for UV-completion through classicalization.
3.1 Massive K-essence
We begin by examining a quadratic potential, specifically a non-tachyonic mass term:
$V(\phi)=V_{m}(\phi)\equiv\frac{m^{2}}{2}\,\phi^{2},$ (3.2)
with $m>0$. A light k-essence boson, characterized by $m\ll\Lambda_{\ast}$, is ’t Hooft natural, as the limit $m\to 0$ restores the shift symmetry of the Lagrangian density, thereby ensuring the radiative stability of the light scalar. As already mentioned in Section 2.1, UV-completion via classicalization renders a little hierarchy $m\ll\Lambda_{\ast}$ a necessary consistency condition for the k-essence model.
Background:
[…]
Linear Regime $(r\gg R_{V})$:
In this regime, the kinetic term dominates over the kinetic self-interaction term. Consequently, the mass term should be compared to the kinetic term:
[…]
Transition Regime $(r\sim R_{V})$:
Here, the kinetic term and the kinetic self-interaction term are of comparable magnitude. The comparison yields:
[…]
leading to the same conclusions as in the linear regime.
Non-linear Regime $(r\ll R_{V})$:
In this regime, the kinetic self-interaction term dominates over the kinetic term. Thus, the mass term should be compared to the former:
[…]
We now compare these scaling arguments with the exact solution for massive k-essence in the presence of a localized source. The Euler-Lagrange equation is provided by Eq. (3.1), with the potential specified as $V(\phi)=V_{m}(\phi)$. In the linear regime, the background solution takes the form
[…]
from Eq. (2.15), with the Heaviside step function $\theta(z)$.
[…]
Perturbations:
For the fluctuation $\delta\phi(x)$ on top of the background $\overline{\phi}(r)$, the quadratic potential introduces a mass term as follows:
$V_{m}(\overline{\phi}+\delta\phi)\supset\frac{m^{2}}{2}\,\delta\phi^{2}=\frac{\overline{m}^{2}}{2}\,\delta\phi_{Z}^{2},\quad\overline{m}(r_{0})=\frac{m}{\sqrt{Z_{\phi}(r_{0})}}.$ (3.8)
Here, we define the renormalized field $\delta\phi_{Z}(x)$ from Eq. (2.28) to identify the effective mass $\overline{m}(r_{0})$ of the fluctuations of the background $\overline{\phi}(r)$. Deep within the Vainshtein core, where $Z_{\phi}(r_{0})\gg 1$, the mass undergoes a redshift, yielding $\overline{m}(r_{0})\ll m$. Consequently, the mass term does not disturb the screening mechanism, and the background solution is perturbatively stable.
Classicalization:
[…]
We have thereby substantiated the conjecture presented in Ref. [108]: UV-completion through classicalization requires at least a little hierarchy, specifically $m\ll\Lambda_{\ast}$. In the absence of such a hierarchy, unitarity should be violated, rendering the theory inconsistent. This outcome exemplifies the concrete impact of UV/IR mixing on the hierarchy of scales.
3.2 K-chameleon: Kinetically Catalyzed Chameleon Screening
We have observed that a quadratic potential for the k-essence field $\phi(x)$, i.e., a mass term, imposes a little hierarchy for Vainshtein screening to take effect. It is therefore instructive to examine more complex potentials. Motivated by SSB, we focus on potentials invariant under the $\mathbb{Z}_{2}$ symmetry $\phi(x)\mapsto-\phi(x)$; however, the same methodology can be applied to study more general potentials.
3.2.1 Symmetric Vacuum
Consider the previously discussed massive k-essence model with a perturbative $\phi^{4}$ interaction, where the potential exhibits a vanishing vacuum expectation value (VEV):
$V(\phi)=V_{\oplus}(\phi)\equiv\frac{m^{2}}{2}\,\phi^{2}+\frac{\lambda}{4!}\,\phi^{4},$ (3.10)
with $m,\ \lambda>0$. In this case, the mass term is not protected by a shift symmetry, rendering a hierarchy $m\ll\Lambda_{\ast}$ unnatural in the ’t Hooft sense, analogous to the Higgs boson mass in the standard electroweak theory [6]. Nonetheless, as our analysis of the quadratic potential suggests, such a modest hierarchy is essential for classicalization to occur.
Vainshtein Screening:
[…]
Linear Regime $(r\gg R_{V})$:
In this regime, the kinetic term dominates over the kinetic self-interaction. To assess the relative importance of the $\phi^{4}$-interaction, we compare it to the kinetic term:
$\left|\dfrac{\lambda\,\phi^{4}}{\partial^{\mu}\phi\partial_{\mu}\phi}\right|\underset{\phi(x)\equiv\overline{\phi}(r)}{\sim}\lambda\left(\frac{R_{V}}{\ell_{\ast}}\right)^{4},$ (3.11)
which is $\ll 1$ when $R_{V}\ll\ell_{\ast}$, i.e., within the standard EFT regime where $\mathcal{E}$ and $\phi(x)\ll\Lambda_{\ast}$. However, even for a moderately weak coupling (e.g., $\lambda\sim 0.1$), the ratio becomes $\gg 1$ when $\mathcal{E}\gg\Lambda_{\ast}$. In this case, the potential dominates over the linear term, and the background solution $\overline{\phi}(r)$ approaches the VEV of $V(\phi)$.
Transition Regime $(r\sim R_{V})$:
Here, the kinetic term and the kinetic self-interaction are of comparable magnitude, yielding:
$\left|\dfrac{\lambda\,\phi^{4}}{\partial^{\mu}\phi\partial_{\mu}\phi}\right|\sim\left|\dfrac{\lambda\,\phi^{4}}{\ell_{\ast}^{4}(\partial^{\mu}\phi\partial_{\mu}\phi)^{2}}\right|\underset{\phi(x)\equiv\overline{\phi}(r)}{\sim}\lambda\left(\frac{R_{V}}{\ell_{\ast}}\right)^{4},$ (3.12)
and thus the conclusions remain consistent with those of the linear regime.
Non-linear Regime $(r\ll R_{V})$:
In this regime, with $\mathcal{E}\gg\Lambda_{\ast}$, the kinetic self-interaction dominates the kinetic term, so we compare the $\phi^{4}$-interaction to the kinetic self-interaction:
[…]
This scaling estimate reveals that a Vainshtein core for $r\gg\ell_{\ast}$ can only exist if $\lambda\ll 1$ to sufficiently suppress the growth of $(r/\ell_{\ast})^{4}\gg 1$. Consequently, for applications involving a moderately weak coupling, a clear tension arises with Vainshtein screening.
Figure 8: Comparison of the scalar potentials $V_{\oplus}(\phi)$ in Eq. (3.10) and $\mathcal{V}_{\oplus}(\phi)$ in Eq. (3.14) with a $\mathbb{Z}_{2}$-symmetric vacuum. The 2 potentials begin to deviate at $|\phi|\sim\Lambda_{\phi}=\Lambda_{\ast}$, the threshold at which chameleon screening of the $\phi^{4}$ interaction becomes effective.
Chameleon Screening:
To resolve this tension between a quartic potential and Vainshtein screening, we seek a minimal modification of the model that satisfies the following 2 criteria: $(i)$ a moderately weak $\phi^{4}$ coupling (e.g., $\lambda\sim 0.1$ or even 1) when $\partial\phi\ll\Lambda_{\ast}^{2}$; $(ii)$ a potential that approaches a mass term as $\partial\phi\gg\Lambda_{\ast}^{2}$. A solution is to recognize that, within the Vainshtein core, both $|\overline{\phi}^{\,\prime}(r)|\gg\Lambda_{\ast}^{2}$ and $|\overline{\phi}(r)|\gg\Lambda_{\ast}$ hold. We therefore modify the potential as follows:
$V(\phi)=\mathcal{V}_{\oplus}(\phi)\equiv\frac{m^{2}}{2}\,\phi^{2}+\frac{\lambda}{4!}\,\Lambda_{\phi}^{4}\tanh\left[\left(\frac{\phi}{\Lambda_{\phi}}\right)^{4}\right],$ (3.14)
where, following Dirac naturalness, we set [^21] A hierarchy $\Lambda_{\phi}\ll\Lambda_{\ast}$ is typically excluded, since the EFT would become strongly coupled for hard scattering processes at energies $\sqrt{s}\sim\Lambda_{\phi}$, well below the classicalization scale $\Lambda_{\ast}$. At such energies, the kinetic self-coupling responsible for classicalization has not yet become operative to unitarize the theory. the new scale $\Lambda_{\phi}\equiv 1/\ell_{\phi}\sim\Lambda_{\ast}$. This example of $V(\phi)$ meets the 2 criteria outlined previously, with the following asymptotic limits (see Fig. 8):
$\Lambda_{\phi}^{4}\tanh\left[\left(\frac{\phi}{\Lambda_{\phi}}\right)^{4}\right]\sim\begin{cases}\phi^{4}&\text{for }|\phi|\ll\Lambda_{\phi},\\ \Lambda_{\phi}^{4}&\text{for }|\phi|\gg\Lambda_{\phi}.\end{cases}$ (3.15)
The second condition guarantees that the mass term governs the potential within the Vainshtein core.
Light scalar fields whose effective potential exhibits strong environmental dependence—commonly referred to as ‘chameleons’—are well-documented in the dark energy literature [226, 227]. In these models, the potential is effectively modified in dense environments, resulting in a large effective mass for the field in regions where its amplitude $|\phi(x)|$ exceeds a critical threshold [^22] Another approach involves achieving an effectively suppressed coupling to matter: see, e.g., the dilaton screening [228, 229] or the ‘symmetron’ mechanism [230, 231, 232]. (a phenomenon known as ‘chameleon screening’). It is feasible to integrate chameleon fields with Vainshtein screening, a hybrid framework referred to as ‘k-chameleons’ [233].
[…]
In the present framework, Vainshtein screening leads to large field values, $|\overline{\phi}(r)|\gg\Lambda_{\phi}$, such that the self-interactions in the scalar field potential are strongly suppressed within the Vainshtein core—a phenomenon we continue to term ‘chameleon screening’. Notably, the region in which this new screening mechanism operates extends beyond the Vainshtein core. Specifically, the departure from a quartic potential becomes significant when $|\overline{\phi}(r)|\sim\Lambda_{\phi}$, which occurs at the ‘chameleon radius’:
$R_{C}\equiv\ell_{\phi}\left(\frac{R_{V}}{\ell_{\ast}}\right)^{2}=\ell_{\phi}\left(\frac{\mathcal{E}}{\Omega\Lambda_{\ast}}\right)\gg R_{V}\gg\ell_{\ast}\sim\ell_{\phi},$ (3.16)
[…]
  • $\spadesuit$
    $R_{V}$ defines the boundary of the Vainshtein core. Within this region, kinetic self-interactions dominate the dynamics;
  • $\spadesuit$
    $R_{C}$ marks the outer limit of the region where chameleon screening takes effect (designated as the ‘chameleon halo’, as it envelops the Vainshtein core);
  • $\spadesuit$
[…]
3.2.2 Tachyon Condensation
The canonical example of the SSB of the $\mathbb{Z}_{2}$ symmetry through tachyon condensation is provided by the following renormalizable potential:
$V_{\ominus}(\phi)$ $\equiv\frac{\lambda}{4!}\left(\phi^{2}-v^{2}\right)^{2},$ (3.17)
$=-\frac{1}{2}\left(\frac{\lambda v^{2}}{6}\right)\phi^{2}+\frac{\lambda}{4!}\,\phi^{4}+\frac{\lambda v^{4}}{24},$ (3.18)
where $\lambda>0$, and $v$ is the VEV of $\phi(x)$. As discussed in Section 3.2.1, a quartic term must be modified to accommodate Vainshtein screening. However, since the quadratic term is tachyonic, it cannot dominate for $|\phi|\gg\Lambda_{\phi}$ without compromising the stability of the system. To address these challenges, we consider the following potential:
$V(\phi)=\mathcal{V}_{\ominus}\equiv\frac{\lambda}{4!}\,\Lambda_{\phi}^{4}\tanh\left[\dfrac{\left(\phi^{2}-v^{2}\right)^{2}}{\Lambda_{\phi}^{4}}\right],$ (3.19)
with the Dirac natural choice $\Lambda_{\phi}\sim\Lambda_{\ast}$. This potential satisfies the asymptotic behavior (see Fig. 10):
$\mathcal{V}_{\ominus}\sim\begin{cases}V_{\ominus}(\phi)&\text{for }|\phi|\ll\Lambda_{\phi},\\ \Lambda_{\phi}^{4}&\text{for }|\phi|\gg\Lambda_{\phi}.\end{cases}$ (3.20)
A notable feature of this model is the restoration of shift symmetry for large field values, $|\phi|\gg\Lambda_{\phi}$, as the potential approaches a constant in this regime.
Figure 10: Comparison of the scalar potentials $V_{\ominus}(\phi)$ in Eq. (3.17) and $\mathcal{V}_{\ominus}(\phi)$ in Eq. (3.19) with tachyon condensation. The 2 potentials begin to deviate at $|\phi|\sim\Lambda_{\phi}=\Lambda_{\ast}$, the threshold at which chameleon screening becomes effective.
[…]
3.2.3 Radiative Stability and Classicalization
Quantum Stability:
The models presented in Sections 3.2.1 and 3.2.2 feature potentials with a specific asymptotic form for $|\phi(x)|\gg\Lambda_{\phi}$:
$\tanh\left[\left(\frac{\phi}{\Lambda_{\phi}}\right)^{4}\right]\underset{|\phi|\gg\Lambda_{\phi}}{\sim}1-2e^{-2\left(\frac{\phi}{\Lambda_{\phi}}\right)^{4}},$ (3.21)
which must remain stable [^23] It is important to note, in general, that fluctuations of the background may exhibit tachyonic instabilities. Instead of ghosts, this is not pathological [3]: classicalon solutions do not correspond to stable field configurations; instead, they decay into a large number of soft quanta. under quantum fluctuations of the background, $\delta\phi(x)=\phi(x)-\overline{\phi}(r)$. To demonstrate this stability, we observe that this asymptotic form can be expressed as
$e^{-2\left(\frac{\overline{\phi}+\delta\phi}{\Lambda_{\phi}}\right)^{4}}=e^{-2\left(\frac{\overline{\phi}}{\Lambda_{\phi}}\right)^{4}}\sum_{n=0}^{N_{0}}\left(\frac{\delta\phi}{\overline{\phi}}\right)^{n}P_{n}\left(\frac{\overline{\phi}}{\Lambda_{\phi}}\right)+o\left[\left(\frac{\delta\phi}{\overline{\phi}}\right)^{N_{0}}\right],$ (3.22)
with polynomial functions $P_{n}(z)$. For $r\ll R_{C}$, the background field satisfies $|\overline{\phi}(r)|\gg\Lambda_{\phi}$, ensuring that all the $\delta\phi(x)$ interaction terms arising from the potential are suppressed by the exponential prefactor. This suppression persists even after field strength renormalization, $\delta\phi(x)\to\delta\phi_{Z}(x)$, as defined in Eq. (2.28).
From this discussion, it is evident [^24] For $|\overline{\phi}|\gg\Lambda_{\phi}$, the exponential factor in Eq. (3.22) always dominates over both polynomial and logarithmic factors of $|\overline{\phi}|/\Lambda_{\phi}$. that radiative corrections cannot alter the form of the potential for $|\phi|\gg\Lambda_{\phi}$ in a manner that would induce vacuum instability (see Ref. [234] for a review on false vacuum decay). This chameleon screening can therefore serve as a stabilization mechanism for the vacuum in theories involving scalar fields [^25] For discussions on vacuum stability in another class of non-local theories, see Refs. [235, 236]..
Classicalization:
We have examined the conditions under which a classicalon solution can form in the presence of a pointlike external source. In a scattering process, the self-interactions of the field serve as the effective source. As demonstrated in Ref. [109], only self-interaction terms involving derivatives give rise to classicalization, in contrast to those arising from a potential. The origin lies in the underlying Vainshtein screening mechanism, which emerges at a substantial distance from the source and thus necessitates derivative interactions [3]. As a result, the classicalization radius $R_{\circledast}$ for the k-chameleon—below which the k-chameleon field loses the ability to resolve smaller length scales—continues to be governed by the kinetic self-interaction and is defined by Eq. (2.35).
Hierarchy Problem:
Upon integrating out the fuzzyons at the scale $\Lambda_{\ast}$, along with the infinite tower of classicalons, one may question the implications for the hierarchy problem with respect to the mass $m$ of the fluctuations around the VEV of the k-chameleon field. Due to the exponential suppression $\sim e^{-N_{\circledast}}$ of an effective vertex involving a classicalon $(N_{\circledast}\gg 1)$ and $\mathcal{O}(1)$ particles—as discussed in Section 2.1—the contributions of classicalons to the threshold corrections $\delta m^{2}$ to $m^{2}$ are likewise exponentially suppressed. The radiative corrections $\delta m^{2}$ are therefore dominated by the heavier fuzzyons—lying at the scale $\Lambda_{\mathrm{fuzz}}$—since $N_{\circledast}\sim 1$ for them. The scale $\Lambda_{\mathrm{fuzz}}$ can be defined as the mass $M_{\circledast}$ for which $R_{\circledast}=\ell_{\ast}$ in Eq. (2.35), yielding $\Lambda_{\mathrm{fuzz}}=\Omega\Lambda_{\ast}$. This relationship leads to the following scaling behavior:
$\delta m^{2}\sim\left(\frac{\Lambda_{\mathrm{fuzz}}}{\Omega}\right)^{2}=\Lambda_{\ast}^{2}.$ (3.23)
From naturalness alone, this suggests that $m\sim\Lambda_{\ast}$. However, classicalization unitarizes hard scattering amplitudes $(\sqrt{s}\gg\Lambda_{\ast})$ by converting the exchange of a few hard bosons (strong coupling) into the production of $N_{\circledast}\gg 1$ soft bosons (weak coupling). Consequently, the k-chameleon bosons must be sufficiently light to be described within the weakly interacting EFT below the scale $\Lambda_{\ast}$, thereby satisfying the little hierarchy $m\ll\Lambda_{\ast}$. This result shows how UV/IR mixing is realized in the language of the $S$-matrix.
3.3 Interaction with Matter
In this section, we address the coupling of a k-chameleon boson to matter, with a focus on its interaction with a fermion field. Given the prevalence of Yukawa couplings in particle physics, our discussion will center on this specific interaction. The objective is to determine how such an interaction with matter can maintain the integrity of Vainshtein screening.
3.3.1 Yukawa Coupling vs Vainshtein Screening
Consider a model featuring 1 real scalar field $\phi(x)$ and 1 Dirac fermion $\Psi(x)$, described by the Lagrangian density $\mathcal{L}_{X}-V(\phi)+\mathcal{L}_{\Psi}$, where
$\mathcal{L}_{\Psi}=\frac{i}{2}\,\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi-y\left[(\Psi_{R}^{\dagger}\cdot\phi)\Psi_{L}+\Psi_{L}^{\dagger}(\phi\cdot\Psi_{R})\right],$ (3.24)
with a real Yukawa coupling $y$. Here, the Dirac matrices $\gamma^{\mu}$ are expressed in a chiral basis. Motivated by a model with a $\mathbb{Z}_{2}$ symmetry (spontaneously broken or not by the potential $V(\phi)$), we assign distinct parities to the Weyl spinors:
$\mathbb{Z}_{2}:\quad\phi(x)\mapsto-\phi(x),\quad\Psi_{L}(x)\mapsto+\Psi_{L}(x),\quad\Psi_{R}(x)\mapsto-\Psi_{R}(x).$ (3.25)
This assignment ensures that the Yukawa interactions in $\mathcal{L}_{\Psi}$ are the only terms invariant under the $\mathbb{Z}_{2}$ symmetry, which introduces an additional contribution to the breaking of shift symmetry, already violated by $V(\phi)$.
By selecting a potential $V(\phi)$ compatible with Vainshtein screening, the background solution $\overline{\phi}(r)$ for $r\ll R_{C}$ is provided by Eq. (2.16), and we add the perturbations $\delta\phi(x)=\phi(x)-\overline{\phi}(r)$. The fermionic Lagrangian density is then modified to
$\mathcal{L}_{\Psi}=\frac{i}{2}\,\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi-\overline{M_{\Psi}}(r)\,\delta\phi_{Z}\overline{\Psi}\Psi-\overline{y}(r)\,\overline{\Psi}\Psi,$ (3.26)
where $\delta\phi_{Z}$ is the renormalized field defined in Eq. (2.25), and
$\overline{M_{\Psi}}(r)=y\,\overline{\phi}(r),\quad\overline{y}(r)=\frac{y}{\sqrt{Z_{\phi}(r)}}.$ (3.27)
As a result, the fermion acquires an effective mass $\overline{M_{\Psi}}(r)$ within the classicalon, which increases as one approaches the source $(r\to 0)$, while the effective Yukawa coupling $\overline{y}(r)$ decreases. A natural question arises regarding the impact of the Yukawa coupling on Vainshtein screening. This issue becomes particularly clear when examining the threshold corrections to the mass of the perturbations $\delta\phi_{Z}(x)$—after field strength renormalization in Eq. (2.28)—through fermion loop corrections inside the Vainshtein core:
$\delta\overline{m}^{2}\sim\left(\dfrac{\overline{y}}{\Omega}\,\overline{M_{\Psi}}\right)^{2}\quad\Longrightarrow\quad\frac{\delta\overline{m}^{2}(r)}{\Lambda_{\ast}^{2}}\sim\frac{y^{4}}{\Omega^{2}}\left(\frac{r}{\ell_{\ast}}\right)^{2}\gg 1,$ (3.28)
for a moderately weak Yukawa coupling, and $R_{V}\gg r\gg\ell_{\ast}$. The fluctuations consequently receive substantial quantum corrections from the Yukawa coupling, which destabilizes the classicalon solution. The issue becomes even more pronounced within the chameleon halo, where $Z_{\phi}\sim 1$. Thus, the model requires modification.
3.3.2 Conformal Coupling and Chameleon Screening
The first modification to the model of Section 3.3.1 introduces a conformal coupling between the k-chameleon and the fermion [^26] For a conformal coupling in the context of chameleon dark energy, see, e.g., Refs. [226, 227].. The Lagrangian density $\mathcal{L}_{\Psi}$ is accordingly modified to [^27] For simplicity, we adopt an identical conformal coupling for both chiralities.
$\mathcal{L}_{\Psi}=\frac{i}{2}\,e^{\left(\frac{\phi}{\Lambda_{\Psi}}\right)^{2}}\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi-y\,\phi\overline{\Psi}\Psi,$ (3.29)
with the scale $\Lambda_{\Psi}\equiv 1/\ell_{\Psi}\sim\Lambda_{\ast}$, which is Dirac natural. A hierarchy $\Lambda_{\Psi}\ll\Lambda_{\ast}$ poses an issue for unitarity, as classicalization is triggered at the scale $\Lambda_{\ast}$.
To elucidate the effect of the conformal factor on the Yukawa coupling, we perform the field redefinition:
$\Psi(x)\mapsto e^{-\frac{1}{2}\left(\frac{\phi}{\Lambda_{\Psi}}\right)^{2}}\Psi(x).$ (3.30)
Under this transformation, $\mathcal{L}_{\Psi}$ becomes [^28] Note that if $\Psi(x)$ transforms non-trivially under a gauge group, the partial derivative $\partial_{\mu}$ is replaced by a covariant derivative without modifying the discussion, as the gauge connection commutes with the conformal factor.
$\mathcal{L}_{\Psi}=\frac{i}{2}\,\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi-y_{C}[\phi]\,\phi\overline{\Psi}\Psi,$ (3.31)
where the Yukawa coupling acquires an effective $\phi$-dependence given by
$y_{C}[\phi]\equiv y\,e^{-\left(\frac{\phi}{\Lambda_{\Psi}}\right)^{2}}.$ (3.32)
Consequently, a background with $|\phi|\gg\Lambda_{\Psi}$ leads to a suppressed coupling $y_{C}[\phi]$, reflecting another manifestation of the underlying chameleon screening mechanism. Like $V(\phi)$, this operator serves as another example of a term that breaks the shift symmetry—otherwise preserved by the UV-screener responsible for classicalization—and hinders Vainshtein screening, thereby requiring suppression via chameleon screening.
By analyzing the background fluctuations $\delta\phi(x)\equiv\phi(x)-\overline{\phi}(r)$ using the classicalon solution (2.16), we obtain the action (3.26) (retaining only the linear term in $\delta\phi$), but with the substitution $y\mapsto y_{C}[\overline{\phi}(r)]$ in the effective parameters defined in (3.27). Within the Vainshtein core ($r\ll R_{V}$), the radiative corrections in Eq. (3.28) are then modified to
$\frac{\delta\overline{m}^{2}(r)}{\Lambda_{\ast}^{2}}\sim\frac{y^{4}}{\Omega^{2}}\,e^{-4\left[\frac{\overline{\phi}(r)}{\Lambda_{\Psi}}\right]^{2}}\left(\frac{r}{\ell_{\ast}}\right)^{2}\ll 1.$ (3.33)
Notably, this exponential suppression extends to the chameleon halo, where $Z_{\phi}\sim 1$, since chameleon screening dominates over Vainshtein screening in suppressing the Yukawa coupling. Regarding the non-linear term in $\delta\phi$ arising from the Taylor expansion of the conformal factor within $y_{C}[\phi]$, the resulting formula closely mirrors that of Eq. (3.22). This demonstrates that all couplings inherit the exponential suppression both inside the Vainshtein core and the chameleon halo. Thus, the classicalon solution remains protected from quantum corrections arising from the fermion coupled to the k-chameleon.
3.3.3 Kinetic Coupling and Classicalization
A persistent challenge arises when probing distances smaller than the scale $\ell_{\Psi}\sim\ell_{\ast}$ by localizing a fermion wavepacket constructed from the field $\Psi(x)$. Since the Yukawa terms, modulated by the conformal factor, lack derivative interactions of both fields, they fail to initiate classicalization at the scale $\Lambda_{\Psi}\sim\Lambda_{\ast}$ and thus cannot self-complete the theory. The second modification to the initial Yukawa theory of Section 3.3.1 is therefore to introduce a kinetic coupling between $\Psi(x)$ and $\phi(x)$:
$\mathcal{L}_{\Psi}=\frac{i}{2}\,\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi-y\,e^{-\left(\frac{\phi}{\Lambda_{\Psi}}\right)^{2}}\phi\overline{\Psi}\Psi+c_{2}^{\prime}\left(\frac{\partial^{\mu}\phi\partial_{\mu}\phi}{2\Lambda_{\ast}^{2}}\right)\left(\frac{i\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi}{2\Lambda_{\ast}^{\prime\,2}}\right),$ (3.34)
with a new scale $\Lambda_{\ast}^{\prime}\sim\Lambda_{\Psi}\sim\Lambda_{\ast}$ and $c_{2}^{\prime}\equiv\pm 1$. Note that the new term preserves the shift symmetry of the original k-essence model.
Pointlike Source:
We reconsider the scenario in which $\phi(x)$ is directly coupled to an external pointlike source. Evidently, the final term of the Lagrangian density (3.34) is dominated by the kinetic self-coupling of $\phi(x)$ within the Vainshtein core [^29] Outside the Vainshtein core, the $\phi$$\Psi$ kinetic interaction term is suppressed relative to the kinetic term of $\phi(x)$ by $\Lambda_{\ast}^{2}\Lambda_{\ast}^{\prime\,2}$., analogous to the kinetic term of $\phi(x)$ itself. Consequently, the $\phi$$\Psi$ kinetic interaction is treated perturbatively. Following the procedure outlined in Section 2.2.3, we focus on a point $\mathbf{r_{0}}$ inside the classicalon but distant from the source. The effective kinetic term of $\Psi(x)$ then becomes
$\frac{iZ_{\Psi}}{2}\,\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi\quad\Longrightarrow\quad\Psi_{Z}(x)\equiv\sqrt{Z_{\Psi}(r_{0})}\,\Psi(x),$ (3.35)
where, inside the Vainshtein core,
$Z_{\Psi}(r)\equiv 1-c_{2}^{\prime}\left[\frac{\overline{\phi}^{\prime}(r)}{\sqrt{2}\Lambda_{\ast}\Lambda_{\ast}^{\prime}}\right]^{2}\underset{|\overline{\phi}^{\prime}|\gg\Lambda_{\ast}^{2}}{\sim}-c_{2}^{\prime}\left[\frac{\overline{\phi}^{\prime}(r)}{\sqrt{2}\Lambda_{\ast}\Lambda_{\ast}^{\prime}}\right]^{2}\quad\Longrightarrow\quad c_{2}^{\prime}=-1.$ (3.36)
Here, we have performed the field strength renormalization of $\Psi(x)$, as its kinetic term is renormalized by the background field $\overline{\phi}(r)$ in Eq. (2.16).
The Yukawa coupling modulated by the conformal factor—second term of Eq. (3.34)—is treated as described in Section 3.3.2, but with the definition of $\overline{y}(r)$ in Eq. (3.27) replaced by
$\overline{y}(r)=\frac{y_{C}\left[\overline{\phi}(r)\right]}{Z_{\Psi}\sqrt{Z_{\phi}(r)}},$ (3.37)
with the same qualitative physical conclusions.
The $\phi$$\Psi$ kinetic interaction term introduces couplings between the background fluctuations $\delta\phi(x)$ and $\Psi(x)$. This yields the following cubic and quartic interactions:
$\partial_{r}\delta\phi_{Z}\left(\frac{i\overline{\Psi_{Z}}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi_{Z}}{2\overline{\Lambda_{\ast}^{\prime}}^{2}}\right)-\left(\frac{\partial^{\mu}\delta\phi_{Z}\partial_{\mu}\delta\phi_{Z}}{2\overline{\Lambda_{\ast}}^{2}}\right)\left(\frac{i\overline{\Psi_{Z}}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi_{Z}}{2\overline{\Lambda_{\ast}^{\prime}}^{2}}\right),$ (3.38)
where the blueshifted interaction scales are defined as
$\overline{\Lambda_{\ast}}\equiv\sqrt{Z_{\phi}(r_{0})}\,\Lambda_{\ast}\gg\Lambda_{\ast},\quad\overline{\Lambda_{\ast}^{\prime}}\equiv\sqrt{Z_{\Psi}(r_{0})}\,\Lambda_{\ast}^{\prime}\gg\Lambda_{\ast}^{\prime}.$ (3.39)
Thus, the quantum version of Vainshtein screening extends to the $\phi$$\Psi$ kinetic interaction, and the corresponding radiative corrections are suppressed accordingly inside the Vainshtein core.
Classicalization:
Regarding the non-perturbative unitarization of hard scattering amplitudes with $\sqrt{s}\gg\Lambda_{\ast}\sim\Lambda_{\ast}^{\prime}\sim\Lambda_{\Psi}$ through self-sourcing terms, the fermion field $\Psi(x)$ now contributes also to the classicalization dynamics via the $\phi$$\Psi$ kinetic coupling. By restoring $\hbar$-units, one can perform a dimensional analysis [^30] The dimensionality of the fermion field is $\Psi]=E^{\frac{1}{2}}L^{-1}$. analogous to that in Section 2.2.4, yielding a classicalization radius for both derivative couplings at the same length scale $R_{\circledast}$, as defined in Eq. (2.35). This scale thus governs the classicalization dynamics of both the $\phi(x)$ and $\Psi(x)$ fields.
Phenomenologically, one anticipates a strongly coupled fuzzyon dynamics between the 2 fields at the shared scale $\Lambda_{\ast}\sim\Lambda_{\ast}^{\prime}\sim\Lambda_{\Psi}$, while the semi-classical regime is expected to exhibit the evaporation of classicalons into $N_{\circledast}\gg 1$ k-chameleon bosons and fermion-antifermion pairs. A comprehensive description of the classicalization dynamics would necessitate a dedicated study to capture the non-perturbative collective behavior.
4 Conclusion and Outlook
In Section 2, we reviewed the classicalization paradigm using the concrete example of a massless k-essence field. The kinetic self-interaction is widely recognized in the dark energy literature for exhibiting Vainshtein screening around localized sources. Our objective was to reinterpret these features within the framework of self-UV-completion through classicalization and UV/IR mixing. We also summarized the critiques of classicalization in the literature, along with their associated loopholes.
In Section 3, we have explored the conditions under which Vainshtein screening—triggered by the kinetic self-coupling of a gauge singlet scalar—can arise in the presence of either a scalar potential or a Yukawa coupling to matter. These conditions are critical for the theory to achieve self-UV-completion via classicalization. In addition to the well-established violation of the standard positivity bound for the coefficient of the kinetic self-interaction term, we have shown that a chameleon screening mechanism must be integrated within the classicalon. This ensures that Vainshtein screening remains robust against perturbations from the potential and the coupling to matter.
This framework offers a fresh perspective on the existence of light scalar bosons, when they appear to suffer from a little hierarchy problem: the UV/IR mixing intrinsic to the classicalization phenomenon requires such a hierarchy to facilitate a consistent self-UV-completion of the theory. Nevertheless, it is important to emphasize that the specific choices of UV-screeners—including the scalar potentials and conformal couplings—serve merely as illustrative examples of the underlying screening mechanisms. Exploring alternative possibilities in this direction therefore remains a valuable endeavor.
Several avenues exist for future research in multiple directions, and we outline here a non-exhaustive list of possibilities. From a theoretical standpoint, potential extensions of this work include:
  • $\spadesuit$
    An investigation of the perturbative renormalization of the theory in the presence of a classicalon background.
  • $\spadesuit$
    A generalization of our analysis to other UV-screening mechanisms, such as those found in Galileon theories [143, 144, 145, 146] or ‘D-BIon’ fields [237].
  • $\spadesuit$
    The development of a theoretical framework to explore the strongly coupled fuzzyon regime near the scale $\Lambda_{\ast}$.
Additionally, classicalization offers several promising applications for concrete model-building in phenomenology, such as:
  • $\spadesuit$
    Screening mechanisms were originally proposed to conceal new long-range forces in scalar-tensor theories of modified gravity [3, 224]; thus, exploring the implications of classicalization in this context represents a natural extension of this work.
  • $\spadesuit$
    Extending our work to scalar fields that transform non-trivially under gauge groups would allow the Higgs mechanism to be embedded within a classicalization framework. This approach could offer an elegant resolution to hierarchy problems, such as those encountered in the standard electroweak theory [6] and in dark Higgs sectors [2].
  • $\spadesuit$
    The synergy between Vainshtein and chameleon screenings provide a highly effective stabilization mechanism for the shape of a classical potential under quantum corrections. This capability could preserve the plateau of the inflaton potential [4] within an EFT of inflation [162, 166, 167]. Additionally, the potential metastability of the Higgs potential [234] presents another compelling application for this mechanism.
As a final comment, we emphasize that classicalization at a scale $\Lambda_{\ast}\ll\Lambda_{G}$ is not expected to be embeddable within string theory (with the gravitational scale [^31] In scenarios involving extra spatial dimensions, the gravitational scale $\Lambda_{G}$ can be significantly lower than the Planck scale $\Lambda_{P}$ (see, e.g., Ref. [171]). $\Lambda_{G}\lesssim\Lambda_{P}$). Rather, it should be regarded as a hypothetical alternative framework in which a QFT may admit multiple classicalization scales, with $\Lambda_{G}$ serving as the ultimate such scale. While string theory is anticipated to exhibit asymptotic darkness—classicalizing above the scale $\Lambda_{G}$ with a spectrum of black hole states [39]—it is also expected to function as a Wilsonian UV-completion between the string scale and $\Lambda_{G}$ [108, 238]. Thus, string theory represents a hybrid UV-completion, combining elements of both the Wilsonian approach and self-UV-completion, yet it features only a single classicalization scale: $\Lambda_{G}$.
Acknowledgments
The author is deeply grateful to R.T. D’Agnolo, L. Darmé, A. Deandrea, A.M. Iyer, and F.N. Mahmoudi for the stimulating discussions on various aspects of this work. Special thanks are extended to B. Bellazzini and G. Isabella for the insightful exchanges regarding positivity bounds, and to F. Brümmer and G. Cacciapaglia for drawing attention to the parallels between classicalization and Higgsplosion. The author thanks T. Berthelot for his work during the internship.